### v2

# Towards Coordinate- and Dimension-Agnostic Machine Learning for Partial Differential Equations

Abstract
Machine learning methods for data-driven identification of partial differential equations (PDEs) are typically defined for a fixed number of spatial dimensions and a particular choice of coordinates in which the data have been collected. This dependence prevents the learned equation from generalizing to other spaces. In this work, we reformulate the problem in terms of coordinate- and dimension-independent representations, paving the way toward what we might call “spatially liberated" PDE learning. In this work, we propose an approach to learning differential equations by expressing them in a way that is independent of the coordinate system and even the underlying manifold where the equation is defined. This allows us to learn a differential equation in low-dimensional spaces and generalize to higher-dimensional spaces with different geometric properties. We provide extensive numerical experiments that demonstrate that our approach allows for seamless transitions across various geometric contexts. We show that the dynamics learned in one space can be used to make accurate predictions in other spaces with different dimensions, coordinate systems, boundary conditions, and curvatures.
Preprint — October 2025
1 Introduction
Machine learning has traditionally been used to identify and model partial differential equations (PDEs) from observational data (Krischer et al., 1993; González-García et al., 1998; Blechschmidt & Ernst, 2021; Cuomo et al., 2022). In particular, deep learning techniques such as physics-informed machine learning (Karniadakis et al., 2021; Raissi et al., 2019) and operator learning (Lu et al., 2021; Kovachki et al., 2023; Fabiani et al., 2025), often in combination with symbolic regression techniques (Brunton et al., 2016; Cranmer, 2023), can extract spatiotemporal patterns from experimental measurements and (approximately) infer the underlying physical laws. This has been especially useful for describing complex dynamical systems in various scientific and engineering domains, where conventional numerical methods can be computationally expensive (Kochkov et al., 2021), or where we only have access to an incomplete description of the system (Martin-Linares et al., 2023). As a result, learning-based PDE models are used on a wide range of scales, from fine-grained descriptions at the particle level (Lee et al., 2023) to coarse-grained representations in terms of density fields (Psarellis et al., 2024), providing a data-driven alternative to traditional physics-based derivations.
Real-world data are often expressed in domains and coordinate systems that are determined by the experimental acquisition process. For instance, bacterial chemotaxis can be captured as waves in effectively one-dimensional channels (Adler, 1966), or visualized as two-dimensional images — top-down, with upright (Phan et al., 2020) or bottom-up, with inverted microscopy (Morris et al., 2017). All three experimental settings produce data of the same physical process supported in different geometries and coordinate systems. The conventional machine learning approaches for PDE identification are inherently tied to a predefined spatial framework, preventing them from generalizing the learned physical mechanism to different coordinate systems, geometries, or dimensions. The recent work of Psarellis et al. (2024) has proposed a coordinate-independent PDE learning approach using exterior calculus. Their method allows them to learn a time evolution scalar PDE independently of the coordinate system, and they show that the learning generalizes to different coordinate system representations of the same domain.
In this work we push these ideas substantially further, and show that we can learn PDEs not just in a coordinate-independent way but also independently of the data domain, dimension, and geometry.
We propose a framework that allows us to learn a dynamical system (a physical mechanistic model) from data in one domain, and then make accurate predictions in another, when the underlying physics is the same.
We are able to do this by combining ideas from the coordinate-free learning of time evolution scalar PDEs from Psarellis et al. (2024), and by recent work in any-dimensional learning that proposes a theoretical framework for machine learning models that are defined on a fixed set of parameters but can be trained and evaluated in spaces of arbitrary size (Levin & Díaz, 2024; Levin et al., 2025).
The most widely used any-dimensional models are arguably graph neural networks (GNNs), where a model defined on a fixed set of parameters can be trained and evaluated on graphs of any size (and GNNs are completely agnostic to the graph sizes). In order to have a sufficiently expressive machine learning model with a fixed number of parameters while the dimension of the input space grows arbitrarily, one typically relies on some form of symmetry constraint (this is mathematically formalized via representation stability (Farb, 2014)). For example, graph neural networks are equivariant with respect to node permutations.
In this work we learn PDE operators that are invariant with respect to local orthogonal symmetries (Weyl, 2015; Frankel, 2011; Nakahara, 2018; Abraham et al., 2012; Villar et al., 2023), which is a feature observed in many natural phenomena. For these systems, we show how machine learning can achieve “full spatial liberation”, enabling training in any space and coordinates, yet being able to make accurate predictions across all others, regardless of coordinates, but also dimensions, boundary conditions, and spatial intrinsic curvatures (i.e. Riemann curvature tensors (Do Carmo & Flaherty Francis, 1992)). Fusion of training data obtained at different coordinates/geometries is also enabled.
In our experiments, we use neural networks to identify time evolution PDEs of coupled scalar fields, general scalar PDEs, and vector PDEs in a coordinate-free and dimension-free way. In particular, we apply our models to the FitzHugh-Nagumo (Bär et al., 2003) and Barkley (Barkley, 1991) reaction-diffusion model operators; and a modified Patlak-Keller-Segel model, derived from in-situ observations of chemotactic bacteria (Phan et al., 2020). We also illustrate our method on the Helmholtz equation, and the incompressible Navier-Stokes equation. Most of our models are trained on data generated in a periodic one-dimensional space using Cartesian coordinates, and we then demonstrate their ability to generalize to higher-dimensional spaces. These include two- and three-dimensional Euclidean spaces in Cartesian or non-Cartesian coordinates, as well as a two-dimensional submanifold of a sphere in Euclidean space, described using geographic as well as stereographic coordinates. We also consider different boundary conditions in these new spaces, such as a no-flux Neumann boundary.
2 Coordinate-free Operator Learning Across Dimensions
Many differential equations can be formulated in a coordinate-free and domain-free mathematical expression. For example, the heat equation written as $\frac{\partial u}{\partial t}=\sum_{i=1}^{n}\frac{\delta^{2}u}{\delta x_{i}^{2}}$ is not coordinate free, since it is expressed in terms of local coordinates. However the heat equation expressed as $\frac{\partial u}{\partial t}=\Delta u$ (where $\Delta$ is the Laplacian operator of the Riemannian manifold $\mathcal{M}$ where $u$ is defined) is not only coordinate-free, since the Laplacian $\Delta$ is an intrinsic operator of $\mathcal{M}$, but also it is domain-free, since the Laplacian exists for every Riemannian manifold. Therefore, the equation is well defined and can be evaluated for manifolds of any dimension. This procedure can be formalized in terms of differential forms and exterior differential systems (see Appendix A). In this paper we argue that we can learn the PDE in one domain and transfer its learning to other manifolds of different dimensions, as long as the differential equation is defined as a “dimension-agnostic” expression.
PDEs coming from mathematical physics and other fields of science and engineering often can be expressed in a coordinate-free manner in terms of their independent variables and their (unknown) fields. Notable examples are the Navier-Stokes (Wilson, 2011) and, more generally, the equations of continuum mechanics (Marsden & Hughes, 1994), the Fokker-Planck equation (Masoliver et al., 1987), the Kuramoto-Sivashinsky equation (Frankel & Sivashinsky, 1987), and many others.
In this paper we consider systems of scalar and vector PDEs. All of them governed by equations that are invariant under local orthogonal transformations (rotations and reflections) and under diffeomorphic translations (Frankel, 2011; Nakahara, 2018; Abraham et al., 2012). Such systems arise frequently in nature, as microscopic physical interactions typically exhibit these symmetries; thus, the emergent PDEs inherit them at macroscopic scales. We consider the PDE to involve at most second-order [^1] Higher-order will be considered in future works by extending the invariant features discussed in Eq. (1). spatial-derivatives, in which we adopt the power-counting scheme of effective field theory (Kaplan, 2005; Buchalla et al., 2014) with the total differential order of each term defined as the sum of its derivative orders. [^2] This is the power-counting scheme for wave numbers, which is of critical importance when creating an effective field theory for coarse-grained dynamics. Since each spatial derivative introduces a factor of the wave number (e.g. in Fourier space), higher-order terms are increasingly suppressed at low wave numbers. By explicitly tracking the total derivative order, this method provides a more accurate characterization of the dominant large-scale emergent behavior of the system. We then observe that when we apply these restrictions, the resulting class of PDEs turn out to be domain-free and dimension-free and we can therefore generalize the learning to other domains in other dimensions.
Our approach can be summarized in two main steps (see Figure 1):
  1. 1.
    Given data satisfying a partial differential equation, first learn a spatially coordinate-free representation of the PDE.
  2. 2.
    Given a new set of initial and/or boundary conditions in a possibly different coordinate system, or even dimension, use a coordinate-free solver of the PDE to evolve the dynamics.
Coordinate-free integrators are a classical subject of study in numerical analysis (Dinesh et al., 2000).
Figure 1: We aim to demonstrate that it is possible to identify a class of evolutionary PDE systems in a coordinate- and dimension-free manner using neural networks. We package the input in a coordinate- and dimension-free manner. (A) We learn the one-dimensional space right-hand-side operator of the PDEs in a Cartesian coordinate. (B) We can then simulate the dynamics in arbitrary coordinates and dimensions (the figure illustrates the case of two-dimensional curved space).
Generators of invariant fields. Given $N$ scalar fields $\quantity{\Psi_{1},\ldots,\Psi_{N}}$ (i.e. 0-forms on a Riemannian manifold $\mathcal{M}$) the set of all possible orthogonally symmetric operators from 0-forms to 0-forms involving up to second-order derivatives can be obtained from [^3] This follows from the first fundamental theorem of classical invariant theory for the orthogonal group (Weyl, 1946; Spivak, 1999), which has been extensively used in machine learning (Villar et al., 2021; Gregory et al., 2024). All SO($n$)-scalars can be built locally from derivatives of scalars, the metric (Kronecker delta), and the Levi-Civita symbol via basic arithmetic operations and contractions (Spivak, 1999); for O($n$)-invariance the Levi-Civita symbol is excluded. Our claim comes from applying this theorem to the second-order jet space, in which in this work we restrict to scalar expressions with total order of derivatives at most two. For higher-order, we can list all possible features constructed from the metric, the fields, and its derivatives — up to the desired order — to be used as invariant features. :
  • The fields $\Psi_{j}$ where $j=1,\ldots,N$.
  • The Laplacians $\Delta\Psi_{j}$, where $j=1,\ldots,N$.
  • The inner products of the gradients $\langle\gradient\Psi_{j},\gradient\Psi_{k}\rangle$, where $j,k=1,\ldots,N$. Since these are symmetric, there are $N(N+1)/2$ distinct products.
The Laplacians and the inner products of the gradients above are defined with respect to the Riemannian metric of $\mathcal{M}$, and can be expressed in coordinate-free form using the metric structure of the manifold (see Appendix A).
This yields a basis of invariant features $\mathcal{B}(\Psi_{1},\ldots,\Psi_{N})$ (or $\mathcal{B}$ when the input is clear by context),
$\mathcal{B}(\Psi_{1},\ldots,\Psi_{N})=([\Psi_{j}]_{j=1}^{N},\,[\Delta\Psi_{j}]_{j=1}^{N},\,[\langle\nabla\Psi_{j},\nabla\Psi_{k}\rangle]_{1\leq j\leq k\leq N})=:\left(b_{1},\ldots,b_{\frac{N(N+5)}{2}}\right).$ (1)
Coordinate-free learning of evolution PDEs. Given $\Psi_{1}$,…, $\Psi_{N}$ as above we consider the PDE identification problem
$\frac{\partial\Psi_{j}}{\partial t}=\mathcal{F}_{j}(\Psi_{1},\ldots,\Psi_{N},t),\quad j=1,\ldots,N$ (2)
where we assume the $\mathcal{F}_{j}$’s are (possibly non-linear) orthogonally symmetric second order operators defining the right-hand-side of the evolutionary PDE. We consider neural networks that learn $\hat{\mathcal{F}}_{j}(\mathcal{B})$ as functions of the features (1), such that the integration of the corresponding PDE via a coordinate-free integrator approximates the training data. A coordinate-free integrator in this context is one that performs the integration of the PDE in terms of the features $b_{i}$. Note that this approach can be seen as a form of canonicalization (see Kaba et al. (2023); Olver (2003)), where the input data is transformed into a set of canonical invariants (here $\mathcal{B}$) to perform the learning.
Coordinate-free learning of general scalar PDEs. The time evolution PDEs we consider here can be expressed as $\frac{\partial u}{\partial t}=\mathcal{F}(\mathcal{B}(u))$ where $\frac{\partial u}{\partial t}$ is the independent variable. A general PDE is an implicit algebraic constraint. We therefore do not a priori know which are the independent or dependent variables. Instead, we can utilize manifold learning methods, linear (e.g. PCA) for linear PDEs, nonlinear (e.g. autoencoders or kernel-PCA) for nonlinear PDEs, to discover the implicit constraint within the coordinate-free basis. In this way, the PDE operator is represented by a latent model constraining the relationship between the variables in $\mathcal{B}$. After this is learned, it can be ‘extracted’ symbolically using symbolic regression, or used implicitly by numerical solvers. Of course, there is the caveat that, when the PDE is nonlinear, the constraint learned is only guaranteed to be accurate in the region adequately sampled by the training data.
The workflow for a general PDE is as follows:
  1. 1.
    Project the data satisfying the PDE, $u$, to the coordinate- and dimension-free basis $\mathcal{B}(u)$.
  2. 2.
    Use manifold-learning to parametrize the submanifold defined by the data in terms of $\mathcal{B}(u)$.
  3. 3.
    Use the trained model as a representation of the operator on different domains, or extract a symbolic representation of the operator using symbolic regression, before transferring to another domain or dimension.
We implement this method for the Helmholtz equation in Section 3.2. We remark that, in order to solve a higher-dimensional problem, we must have information about the ‘well-posedness’ of that problem. For example, the harmonic oscillator needs only two scalar initial/boundary conditions to be solved, whereas the Helmholtz equation needs continua of boundary conditions to be well-posed. This information is not encoded in our coordinate-free representation of the operator in general. However, for time evolution problems, the required initial/boundary conditions are the same across other domains and dimensions.
Extension to vector PDEs. Our proposed approach can be extended to vector PDEs, or more generally coupled scalar and vector field PDEs, by computing an extended coordinate-free feature basis $\tilde{\mathcal{B}}$, which includes scalar features $\mathcal{S}$ and vector features $\mathcal{V}$. Given this basis, a general class of coupled scalar-vector PDEs takes the form:
$\mathcal{F}(\tilde{\mathcal{B}})=\sum_{i=1}^{|\mathcal{V}|}f_{i}\quantity({\quantity{s_{j}}_{j=1}^{|\mathcal{S}|}})v_{i}=0$ (3)
where $f_{i}:\mathcal{S}\to\mathbb{R}$ are arbitrary nonlinear functions. This equation matches the parameterization of equivariant functions to actions by the orthogonal group from Villar et al. (2021). We bound the degree of the differentials in the basis by $2$ in our work; this can, however, be relaxed based on the problem at hand.
The methodology for constructing coordinate-free features from a scalar field $s$ remains the same as in (1). In particular, the scalar features up to order-$2$ will be $\quantity{s,\Delta s}\in\mathcal{S}$, and the sole vector feature will be $\gradient{s}$. For a vector feature $v$ we will have $\expectationvalue{v,v}\in\mathcal{S}$ and $v,\Delta v\in\mathcal{V}$, where the latter operator is the component-wise Laplacian. Finally, we include terms of the form $\expectationvalue{v,\gradient\cdot v}\in\mathcal{V}$, which can be formulated in a cordinate-free manner in terms of the Lie derivative, and terms of the form $\divergence{v}\in\mathcal{S}$, denoting divergence. The remaining elements of the library are then computed using the inner product, divergence, and Lie-derivative operations, all of which have coordinate-free formulations in terms of exterior algebra.
Any-dimensional coordinate-free learning. The Laplacians and inner products are defined as functions of the input scalar fields for any domain –any Riemannian manifold $(\mathcal{M},g)$ of any dimension $n$. Therefore $\mathcal{B}$ (or analogously $\tilde{B}$ in the vector PDE case) is defined via a domain- and dimension-independent description. The learned PDEs are defined as functions of $N(N+5)/2$ features, $(b_{1},\ldots,b_{N(N+5)/2})$, independently of the size of the domain of the $b_{i}$’s (and thus independently of the dimension of the domain of the inputs).
3 Computational Experiments
3.1 Time-Evolution Systems
In this section, we apply the proposed methodology to three time-evolution PDEs: the FitzHugh-Nagumo, the Barkley, and a modified Patlak-Keller-Segel Systems defined below. Each can be expressed in a coordinate-free and dimension-free form (Appendix C), involving two coupled scalar fields (thus $N=2$ in Eq. (2) and Eq. (1)). We present selected computational experiments in the main text, while extensive additional studies are included in the appendices. See Appendices EL for the experimental setup details and quantitative evaluation of the results.
The FitzHugh-Nagumo System. In the FitzHugh-Nagumo system (Bär et al., 2003) two scalar fields participate: the excitation variable $V(x,t)$ and the recovery variable $W(x,t)$. The spatiotemporal dynamics of these fields are governed by the set of two coupled PDEs:
$\begin{split}\partial_{t}V(x,t)&=\vec{\nabla}^{2}V(x,t)+V(x,t)-\frac{1}{3}V^{3}(x,t)-W(x,t)\ ,\\ \partial_{t}W(x,t)&=\varepsilon\left[V(x,t)+\beta-\gamma W(x,t)\right]\ ,\end{split}$ (4)
where we set $(\varepsilon,\beta,\gamma)=(0.02,0,0.5)$. For this choice of parameters, the system generates spiral waves in two-dimensional space and scroll waves in three-dimensional space, whereas such phenomena do not emerge in dimension one.
The Barkley System. The Barkley reaction-diffusion system (Barkley, 1991) models excitable media and oscillatory media. It is often used as a qualitative model in pattern forming systems like the Belousov-Zhabotinsky reaction and other systems that are well described by the interaction of an activator and an inhibitor component. In the simplest case, the Barkley system is described by a set of two coupled PDEs:
$\begin{split}\partial_{t}U(x,t)&=\vec{\nabla}^{2}U(x,t)+\frac{1}{\varepsilon}U(x,t)\left[1-U(x,t)\right]\left[U(x,t)-\frac{V(x,t)+b}{a}\right]\ ,\\ \partial_{t}V(x,t)&=U(x,t)-V(x,t)\ ,\end{split}$ (5)
where we choose the parameters $(\varepsilon,a,b)=(0.02,0.75,0.02)$ so that this system can generate spiral and scroll waves in two- and three-dimensional spaces respectively.
A Modified Patlak-Keller-Segel System. Recent technological advancements have enabled us to directly measure in situ the time-evolution dynamics of bacterial density and the concentration of environmental cues to which the bacteria respond, with high precision (Phan et al., 2024). We can therefore develop much more accurate macroscopic mathematical models to describe the collective behavior of bacterial swarms. We consider the modified Patlak-Keller-Segel model for bacterial chemotaxis as proposed in Phan et al. (2024), supported by direct in situ observations of E. coli (Escherichia coli) density and Asp (Aspartate) chemoattractants. The dynamics are described by a set of two PDEs, which can be can be written in a non-dimensionalized form as follows:
$\begin{split}\partial_{t}A(x,t)&=\vec{\nabla}^{2}A(x,t)-\left[\frac{A(x,t)}{1+A(x,t)}\right]\left[\frac{B(x,t)}{1+B(x,t)}\right]\ \ ,\\ \partial_{t}B(x,t)&=\alpha B(x,t)+\vec{\nabla}^{2}\left[D_{B}B(x,t)\right]\\ &-\vec{\nabla}\left(\chi_{0}\left[\frac{B(x,t)}{1+B(x,t)/B_{h}}\right]\vec{\nabla}\left\{\ln\left[1+\frac{A(x,t)}{K_{i}}\right]\right\}\right)\ \ ,\end{split}$ (6)
where the parameters are estimated from experiments to be $\alpha\approx 1.5\times 10^{-4}$, $D_{B}\approx 0.55$, $\chi_{0}\approx 4.8$, $B_{h}\approx 1.3$, and $K_{i}\approx 0.27$. The field $A(x,t)$ represents the chemical concentration, and the field $B(x,t)$ represents the bacterial density. The effective chemosensitivity is denoted by $\chi[B(t)]$ and the perceived chemical signal $\Phi[A(t)]$ is defined as:
$\chi[B(t)]=\chi_{0}\left[\frac{B(t)}{1+B(t)/B_{h}}\right]\ ,\ \Phi[A(t)]=\ln\left[1+\frac{A(t)}{K_{i}}\right]\ .$ (7)
Method. To study all these systems, we use the following algorithmic workflow:
  1. 1.
    Given training data of the scalar fields $\quantity{\Psi_{i}(x,t)}_{i=0}^{N}$ in one-dimensional Euclidean space, we project the data onto the spatial coordinate-free basis $\mathcal{B}$ in (1) and estimate the time derivatives $\{\partial\hat{\Psi}_{i}/\partial t\}$ using a finite-difference scheme,
  2. 2.
    We train a fully-connected neural network $\hat{\mathcal{F}}_{\text{NN}}$ to predict $\{\partial\hat{\Psi}_{i}/\partial t\}$ from the basis $\mathcal{B}$,
  3. 3.
    Given the trained network $\hat{\mathcal{F}}_{\text{NN}}$, we integrate the corresponding scalar fields (now defined on different choices of domain) given initial conditions.
One-Dimensional Euclidean Space. We first study the generalization of our coordinate-free representation to arbitrary (nonlinear) parametrizations in the spacial domain. In Figure 2, we demonstrate that the learned representation can be used to accurately predict the time evolution for the Patlak-Keller-Segel system under different choices of coordinates. For details, see Appendix D.1.
Two- and Three- Dimensional Euclidean Space. We study the generalization of the learned FitzHugh-Nagumo and Barkley systems in two and three dimensions, represented in Figure 3. Again, we observe agreement between the true and generalized (learned from 1-D) dynamics. We emphasize that the learned operator produces spiral waves, which are a property of solutions in higher dimensions that cannot be inferred from the 1-D solution. For details, see Appendix D.2.
Generalization to different curvatures and topologies. The proposed coordinate-free representation generalizes to domains with different curvature and topology. In particular, we demonstrate that the learned (from 1-D data) operator accurately predicts the time-evolution of the FitzHugh-Nagumo system on a 2-D domain embedded on the surface of a 3-D ball with nontrivial curvature (see Appendix D.3, Figure 6). In this example, we consider both the geographic and stereographic coordinates of the sphere. In addition, we also consider the case where the domain is a flat punctured disk with nontrivial topology. For details, see Appendix D.4.
Figure 2: The Patlak-Keller-Segel field evolution dynamics in a periodic one-dimensional space, with different coordinate systems. The initial field configurations are the same, up to an extrapolation when transitioning between different coordinate systems. (A) The Cartesian and non-Cartesian coordinate systems, using Eq. (25) with different values of $L$ to convert between $x$ and $y$. (B1) The time progression of the fields when integrating Eq. (6) to generate the true dynamics Cartesian coordinate $x$. (B2) The time progression of the fields when using the trained neural network to generate the learned dynamics in the Cartesian coordinate $x$. (C) Transformation of Fig. (B1) into the non-Cartesian coordinate $y$. (D1) The time progression of the fields when using the trained neural network to generate the learned dynamics in the non-Cartesian coordinate $y$. (D2) Transformation of Fig. (D1) into the Cartesian coordinate $x$.
Figure 3: The time evolution dynamics of FitzHugh-Nagumo and Barkley systems in higher-dimensional Euclidean spaces. We study the FitzHugh-Nagumo system on a two-dimensional domain with periodic boundary conditions under various coordinate representations in (A), and investigate the Barkley system in three-dimensional space subject to Neumann (no-flux) boundary conditions in (B). For (A), we present the time evolution of the true dynamics obtained by integrating Eq. (4) and compare it with the dynamics generated using the learned model at four different time frames: (A1) $t=0$, (A2) $t=15$, (A3) $t=30$, (A4) $t=45$. Each time frame consists of four two-dimensional plots: the top row shows the true dynamics, while the bottom row displays the neural network predictions; the left column corresponds to $V$-field and the right column to $W$-field of the FitzHugh-Nagumo system. For (B), we present the time evolution of the true dynamics obtained by integrating Eq. (5) and compare it with the dynamics generated using the learned model at four different time frames: (B1) $t=0$, (B2) $t=1.5$, (B3) $t=3.0$, (B4) $t=4.5$. Every time frame consists of four three-dimensional plots (each illustrated with two $x_{3}$-slices, located at $x_{3}=\pm 35$): the top row shows the true dynamics, while the bottom row displays the neural network predictions; the left column corresponds to $U$-field and the right column to $V$-field of the Barkley system.
3.2 General PDE Systems
The Helmholtz Equation. Given data satisfying the Helmholtz equation $\Delta u=-ku$, for $k=1$, the goal is to learn the PDE from data in dimension 1 (where it reduces to the harmonic oscillator), and generalize to higher dimensions. The coordinate-free features for one scalar field are $\mathcal{B}=\quantity{u,\expectationvalue{\gradient u,\gradient u},\Delta u}$. Since we only have one PDE, $\mathcal{B}(u)$ should span a two-dimensional manifold that we call $\mathcal{L}$, and since the PDE is linear, PCA should be able to identify it. At this point, the operator is implicitly encoded in the weights and components of the fitted PCA model. (Also note that, after computing derivatives, and embedding in $\mathcal{B}$, we discard all information that has to do with which point corresponds to which solution; they all simply lie on a submanifold in that space.)
Now, suppose that we want to transfer this operator to another domain, for example a 2D unit square equipped with the flat metric. We will do this using a Physics-Informed neural network (PINN) (Raissi et al., 2019), solving the PDE implicitly defined by constraints in the manifold $\mathcal{L}$ and boundary conditions prescribed by the user. In the classical setting, for any given domain with an initial condition, the PINN produces a candidate solution $\hat{u}$. The PDE law and boundary conditions are imposed in the loss function as ‘soft’ constraints. To train the PINN using the implicit coordinate-free representation of the operator, encoded by the prior PCA components, we proceed as follows:
  1. 1.
    Sample a set of points $\quantity{x}$ in the domain $\Omega$
  2. 2.
[…]
  3. 3.
[…]
  4. 4.
[…]
  5. 5.
[…]
    $\mathcal{E}_{\text{PDE}}=\norm{\hat{\hat{u}}_{\mathcal{B}}-\mathcal{B}(\hat{u})}$.
The proposed framework is depicted in Fig. 4. The boundary conditions (BC) are imposed in exactly the same manner as in the classical setting, with a PINN training loss taking the form $\mathcal{E}_{\text{PINN}}=\mathcal{E}_{\text{PDE}}+\mathcal{E}_{\text{BC}}$.
In Fig. 14 (Appendix M) we compare solutions to the 2D Helmholtz equation on the unit square with the flat metric, obtained using (i) PINNs and (ii) the projection-based algorithm described above, where the PDE law is implicitly encoded by the fixed PCA model in coordinate-free space. For details, see Appendix M.
Figure 4: Implicit Coordinate-Free Representation: Left: Solutions of the 1-D Helmloltz problem (harmonic oscillator) in the original variables. Middle: Embedding of the solutions to $\mathcal{B}$, with black arrows depicting the two PCA modes obtained from the data. Right: Embedding of one prediction of a randomly initialized PINN in $\mathcal{B}$ (yellow), along with its projection (red) on the previously learned PCA submanifold . The PDE loss tests whether the yellow and green representations agree, i.e. if the PINN solution resides on the learned submanifold.
Remark: While the PDE constraint itself can be encoded as a relation (linear or nonlinear) between the coordinate-free basis elements in any input space, it does not inform us about what the necessary and sufficient initial and boundary conditions are to solve a PDE in a different geometry or dimension. For a relevant discussion in terms of well-posedness, see Bertalan et al. (2025)). This is not the case for time-evolution problems, where knowledge of the initial field is sufficient.
3.3 Vector PDEs
Figure 5: Magnitude of velocity field for simulated flow around cylinder (top) and integrated (learned, coordinate-free) Taylor-Green Vortex dynamics (bottom)
Incompressible Navier Stokes: We apply our extended methodology to the 2D incompressible Navier-Stokes equations. We consider two well-studied settings, that of flow around a cylinder, and the Taylor-Green Vortex (Taylor & Green, 1937), illustrated in Figure 5. In the latter case, an analytical solution is known. The equation takes the form:
$\partialderivative{v}{t}=-\quantity(v\cdot\gradient)v-\frac{1}{\rho}\gradient p+\frac{\eta}{\rho}\Delta v$ (8)
$\gradient\cdot{v}=0,$ (9)
where $v=(v_{1},v_{2})$ is the velocity vector field, $p$ is the pressure, a scalar quantity, $\rho$ is the density which is assumed to be constant, and $\eta$ is the dynamic coefficient of the incompressible fluid, and is also constant. It is easy to see that this is an instance of the general form (3).
We will consider the problem of learning the momentum equation from the flow around a cylinder, and transferring it to the taylor-green vortex setting. Because of the simplified nature of the incompressible equations, we can perform least-squares to identify the coefficients of the vector features (which play the role of the $f_{i}$ in (3)) instead of using a neural network. We note that the compressible Navier-Stokes would require the $f_{i}$ be nonlinear, at least as a function of density. We follow a familiar pattern: We compute the coordinate-free features of the Navier-Stokes solution in the cylinder setting, and perform least-squares to learn $\partialderivative{v}{t}$ as a function of the spatial-derivative features. We then transfer the operator to the Taylor-Green Vortex setting (assuming the pressure is known), and integrate given an initial condition for which the analytical solution is known. In these two examples, the topology of the domain is different. For details, see Appendix N.
4 Discussion
We propose a framework that, for a general class of PDE systems, allows us to learn and represent the equation in a “spatially liberated” manner. The training can be performed using data from simple 1-D numerical simulations or experimental settings, yet it can generalize across dimensions, coordinate systems, and domain curvatures, even when varying boundary conditions. Training can also be performed “fusing" available data from multiple dimensions or geometries. This flexibility makes our approach particularly valuable at the interface between experiments and engineering, as it enables seamless integration of experimental data from different geometry/dimension experiments into predictive models, allowing theoretical frameworks to adapt to complex real-world geometries.
Using ideas from any-dimensional machine learning, in this work, we focus on identification of second-order PDEs that are invariant under the full local symmetry group of Euclidean space (rotations and reflections, and translations) – which often emerge from long-wavelength physical effective field theories (Kaplan, 2005; Buchalla et al., 2014).The next step is to apply this framework to solving PDEs in a dimension agnostic manner that do not explicitly require the use of integrators (here we use coordinate-free integrators). For example, by defining an any-dimensional formulation of neural operators or PINNs that allow us to transfer PDE solutions from one domain to another.
Implicit to any-dimensional generalization is the underlying assumption of what it means to take a model from one space to another. In the case of graph neural networks assumptions that GNNs make to generalize to different graph sizes have been studied extensively in the graph transferability literature (Ruiz et al., 2020; Maskey et al., 2023; Velasco et al., 2024; Levin et al., 2025). For PDEs, the generalization to other domains is done via the intrinsic manifold operators (Laplacians, exterior derivatives, etc), but more work is needed to fully characterize what these methods are doing.
We expect this type of generalization to be broadly beneficial for the machine learning community, as PDEs play a central role not only in modeling of physical systems, but also optimization algorithms, sampling, and diffusion algorithms, which are fundamental to generative AI. A theoretically grounded approach to distilling information that can be applied to systems of arbitrary dimension could significantly boost efficiency for high-dimensional methods.
Appendix A Background on Differential Forms
The natural mathematical formalism to express coordinate-free operators on Riemannian manifolds is given by exterior calculus, which is built on differential forms. In this section, we give the relevant definitions and background.
A differential form is a mathematical object that generalizes scalars, vectors, and volumes, enabling integration and differentiation on manifolds in a coordinate-free and, to some extent, also in a dimension-free manner. A general $k$-form is a combination of $k$-differentials:
$\Omega=\omega_{\mu_{1}\mu_{2}...\mu_{k}}dx^{\mu_{1}}\wedge dx^{\mu_{2}}\wedge...\wedge dx^{\mu_{k}}\ ,$ (10)
where $\omega_{\mu_{1}\mu_{2}...\mu_{k}}$ are asymmetric tensor components, $\wedge$ denotes the wedge product which is anti-commutative, and note that we use the Einstein summation convention unless otherwise specified. On a Riemannian manifold, the contravariant version of these components can be obtained by raising the indices using the inverse metric tensor $g^{\mu\nu}$ (which is the inverse of the symmetric metric tensor $g_{\mu\nu}$, i.e. $g^{\mu\nu}{g_{\mu\rho}}=\delta^{\nu}_{\rho}$, where $\delta$ is the Kronecker delta tensor):
$\omega^{\mu_{1}\mu_{2}...\mu_{k}}=g^{\mu_{1}\nu_{1}}g^{\mu_{2}\nu_{2}}...g^{\mu_{k}\nu_{k}}\omega_{\nu_{1}\nu_{2}...\nu_{k}}\ .$ (11)
The Hodge star is the operator that turns a $k$-form into a (nk)(n-k)-form:
$\star\Omega=\frac{k!}{(n-k)!}\epsilon_{\mu_{1}\mu_{2}...\mu_{d}}\omega^{\mu_{1}\mu_{2}...\mu_{k}}dx^{\mu_{k+1}}\wedge dx^{\mu_{k+2}}\wedge...\wedge dx^{\mu_{n}}\ ,$ (12)
in which $\epsilon_{\mu_{1}\mu_{2}...\mu_{d}}$ is the Levi-Civita tensor (exists on orientable manifolds). The exterior derivative $d\Omega$ is a $(k+1)$-form, which can be calculated with:
$d\Omega=d\omega_{\mu_{1}\mu_{2}...\mu_{k}}\wedge dx^{\mu_{1}}\wedge dx^{\mu_{2}}\wedge...\wedge dx^{\mu_{k-1}}\wedge dx^{\mu_{k}}\ ,$ (13)
where $d\omega_{\mu_{1}\mu_{2}...\mu_{k}}$ corresponds to the differential of the component $\omega_{\mu_{1}\mu_{2}...\mu_{k}}$, i.e.
$d\omega_{\mu_{1}\mu_{2}...\mu_{k}}=\partial_{\mu_{k+1}}\omega_{\mu_{1}\mu_{2}...\mu_{k}}dx^{\mu_{k+1}}\ .$ (14)
In a coordinate-explicit representation, given the metric structure $g$ defined on the manifold $\mathcal{M}$, the Laplacians and the inner products of gradients from Section 2 can be written as:
  • The Laplacians:
    $\Delta\Psi_{j}=\star\ d\star d\Psi_{j}=\left|g\right|^{-1/2}\partial_{\mu}\left(\left|g\right|^{1/2}g^{\mu\nu}\partial_{\nu}\Psi_{j}\right)\ ,$ (15)
    where $j=1,2,3,$... and $\left|g\right|$ is the determinant of the metric tensor $g_{\mu\nu}$. There are $N$ of them in total.
  • The inner products of the gradients:
    $\langle\nabla\Psi_{j},\nabla\Psi_{k}\rangle=\langle d\Psi_{j},d\Psi_{k}\rangle\equiv\star\left(d\Psi_{j}\wedge\star\ d\Psi_{k}\right)=g^{\mu\nu}\partial_{\mu}\Psi_{j}\partial_{\nu}\Psi_{k}\ ,$ (16)
    where $j,k=1,2,3,$... Since these products are symmetric, there are $N(N+1)/2$ distinct inner products in total.
Appendix B Numerical Scheme
To investigate systems of PDEs numerically, we need to specify the differentiation and integration schemes because they govern how the spatio-temporal continuous PDEs are approximated and then solved in a discrete form. The chosen methods influence the accuracy, stability, computational cost, and the appropriateness of the numerical solution.
B.1 Spatial-Differentiation
Here, the continuous spatial-derivatives in the PDEs are replaced by discrete approximations, i.e. central finite differences at the second-order accuracy. For simplicity, we consider the spatial-differentiation, such as $\partial_{x}$, $\partial_{x}^{2}$, of one-dimensional space scalar field, e.g. $M(x,t)$, $N(x,t)$. Generalizing to higher dimensions is straightforward, as each coordinate is treated independently.
For the first spatial-derivative, given that $\Delta x$ is the spacing between adjacent grid points, we use the following formula at each grid points:
$\partial_{x}M(x,t)\overset{\text{num}}{\approx}\frac{M(x+\Delta x)-M(x-\Delta x)}{2\Delta x}\ .$
For the second spatial-derivative, we apply:
$\partial_{x}^{2}M(x,t)\overset{\text{num}}{\approx}\frac{M(x+\Delta x)-2M(x)+M(x-dx)}{\Delta x^{2}}\ .$
Note that $\partial_{x}^{2}M(x,t)\overset{\text{num}}{\approx}\partial_{x}\left[\partial_{x}M(x,t)\right]$ up to a small error of order $\mathcal{O}(\Delta x^{2})$. When dealing with spatial derivatives inside other spatial derivatives, it can be beneficial to split them out rather than applying the derivatives consecutively before using the numerical approximations, e.g.
$\partial_{x}\left[N(x,t)\partial_{x}M(x,t)\right]\rightarrow\partial_{x}N(x,t)\partial_{x}M(x,t)+N(x,t)\partial_{x}^{2}M(x,t)\overset{\text{num}}{\approx}.$..\ . (17)
This splitting helps control error propagation, maintains higher accuracy, and also provides better handling of boundary conditions, preventing error accumulation near boundaries.
When considering a finite domain e.g. $x\in[-L_{x},+L_{x}]$, the boundary conditions at the edges of the domain i.e. $x=\mp L_{x}$ dictate how we should modify Eq. (B.1) and Eq. (B.1) for points lying outside of the domain. For examples:
  • Periodic boundary conditions:
    $M(\mp L_{x}\mp\Delta x)=M(\pm L_{x})\ .$ (18)
  • Neumann (no-flux) boundary conditions:
    $M(\mp L_{x}\mp\Delta x)=M(\mp L_{x})\ .$ (19)
    We can see directly from this identification that $\partial_{x}M(x=\mp L_{x})=0$.
B.2 Temporal-Integration
Here, we use the forward Euler method to discretize the time derivative with spacing $\Delta t$:
$\partial_{t}M(x,t)\overset{\text{num}}{\approx}\frac{M(x,t+\Delta t)-M(x,t)}{\Delta t}\ ,$ (20)
thus:
$\partial_{t}M(x,t)=\mathcal{F}(t)\ \ \Longrightarrow\ \ M(x,t+\Delta t)=M(x,t)+\mathcal{F}(t)\Delta t\ ,$ (21)
where $\mathcal{F}$ is a function that can depend not only on time position $t$, but also on space location, the fields value, and their spatial-derivatives e.g. the right-hand-side of the PDE. To numerically integrate $M(x,t)$, given the initial field value, we use Eq. (21) consecutively every timestep $t\rightarrow t+\Delta t$.
Appendix C Equations in Coordinate-Free Form
Here we give a coordinate-free expression for the PDEs from Section 3.1 using the formalism from Appendix A.
FitzHugh-Nagumo System:
$\begin{split}\partial_{t}V(t)&=\star\ d\star dV(t)+V(t)-\frac{1}{3}V^{3}(t)-W(t)\ ,\\ \partial_{t}W(t)&=\varepsilon\left[V(t)+\beta-\gamma W(t)\right]\ ,\end{split}$ (22)
Barkley System:
$\begin{split}\partial_{t}U(t)&=\star\ d\star dU(t)+\frac{1}{\varepsilon}U(t)\left[1-U(t)\right]\left[U(t)-\frac{V(t)+b}{a}\right]\ ,\\ \partial_{t}V(t)&=U(t)-V(t)\ ,\end{split}$ (23)
Modified Patlak-Keller-Segel System:
$\begin{split}\partial_{t}A(t)&=\star\ d\star dA(t)-\left[\frac{A(t)}{1+A(t)}\right]\left[\frac{B(t)}{1+B(t)}\right]\ \ ,\\ \partial_{t}B(t)&=\alpha B(t)+\star\ d\star d\left[D_{B}B(t)\right]-\star\ d\star d\left\{\chi[B(t)]\Phi[A(t)]\right\}\ ,\end{split}$ (24)
Appendix D Numerical Results for Time-Evolution PDEs
The learning of all PDE systems is conducted on data from a one-dimensional periodic domain $x\in[-L,+L]$. The data and neural network architectures are detailed in Appendix E. All simulations use evenly spaced grids for finite-difference computations in the specified coordinate system, and we verify that the numerical schemes are robust to space-time discretizations (see Appendix I). For notational convenience, we denote the fields obtained by numerically integrating the exact PDE as $\Psi_{i}$, and the predictions generated by the neural network models as $\hat{\Psi}_{i}$. The argument $\eta(\zeta)$ denotes that the computation was performed in $\zeta$-coordinates and has been push-forwarded into $\eta$-coordinates via a coordinate transformation. A uniform grid in $\zeta$-coordinates will in general correspond to a non-uniform grid in $\eta$-coordinates.
All quantitative comparisons between the “true” and learned dynamics are contained in Appendix K.
D.1 One-Dimensional Euclidean Space
In this example, we learn the Patlak-Keller-Segel field evolution dynamics in a periodic one-dimensional space. We train a network with data in 1-D Cartesian coordinates $x$, and we test its performance under a change of coordinates $y$ in the same 1-D domain. Specifically $y$ is related to $x$ by the monotonic but nonlinear function:
$x=f(y)\equiv x(y)=(y/2)\left[1+(y/\lambda)^{2}\right]\ ,$ (25)
as illustrated in Fig.2A. We set $\lambda=L$. The inverse function $y=f^{-1}(x)\equiv y(x)$ is provided in Appendix G.1. The input pre-processing for the neural network (Fig. 1B, in $y$-coordinate) is detailed in Appendix G.2.
We test the trained model on a new initial condition $t=0$ with uniform $A$ and localized peaks in $B$ (see Appendix I.1). The true dynamics, $A(x,t)$ and $B(x,t)$, are obtained by integrating Eq.(6) (Fig.2B1); the learned dynamics, $\hat{A}(x,t)$ and $\hat{B}(x,t)$, are generated by the neural network (Fig.2B2) are a good approximation. To test our coordinate-free generalization claims, we interpolate the initial field configurations to $y$-coordinates (see Appendix H), integrate the neural network model to get $\hat{A}(y,t)$ and $\hat{B}(y,t)$ (Fig.2D1), and then map the results back to $x$-space to arrive at $\hat{A}(x(y),t)$ and $\hat{B}(x(y),t)$ (Fig.2D2). We can also transform Fig. 1B1 to $y$-coordinates, obtaining $A(y(x),t)$ and $B(y(x),t)$ as shown in Fig. 2C. We observe that Fig. 2C and Fig. 2D1 agree, as well as Fig. 2A1 and Fig. 2D2, demonstrating that our machine learning method generalizes across coordinate systems.
D.2 Two- and Three-Dimensional Euclidean Space
In this section we consider the FitzHugh-Nagumo system and the Barkley system trained in 1-D cartesian data. We first add spatial-dimension to the FitzHugh-Nagumo system during test time. Using the trained models, we predict dynamics in both Cartesian $(x_{1},x_{2})$ and non-Cartesian $(y_{1},y_{2})$ coordinates, related by:
$x_{1}=f(y_{1})\equiv x_{1}(y_{1})\ ,\ x_{2}=f(y_{2})\equiv x_{2}(y_{2})\ ,$ (26)
where $f$ is as in Eq.(25) and we set $\lambda=L$. The domain is the square $(x_{1},x_{2})\in[-L,+L]\times[-L,+L]$ with periodic boundaries. An illustration of $(y_{1},y_{2})$ versus $(x_{1},x_{2})$ is in Appendix G.1; input pre-processing in $(y_{1},y_{2})$-coordinates for the neural network is detailed in Appendix G.3.
Spiral waves can emerge for certain instances of the FitzHugh-Nagumo system. We choose an initial field configuration at $t=0$ where the spiral wave pattern is clearly visible (an explanation of how we do it is provided in Appendix I.2). In $(x_{1},x_{2})$-coordinates, the true dynamics are computed by integrating Eq.(4) (Fig. 3A1-4 top row, $V(x_{1},x_{2},t)$ and $W(x_{1},x_{2},t)$). In $(y_{1},y_{2})$-coordinates, the learned model is integrated from the extrapolated initial condition (see Appendix H for details), and the results are mapped back to $(x_{1},x_{2})$ (Fig. 3A1-4 bottom row, $\hat{V}(x_{1}(y_{1}),x_{2}(y_{2}),t)$ and $\hat{W}(x_{1}(y_{1}),x_{2}(y_{2}),t)$). We observe that the predictions agree, suggesting that our machine learning method not only generalizes across coordinate systems, but also it extends across dimensions. The correct predictions are made in two dimensions, even though the model was trained in one dimension, indicating that the approach is dimension-free.
We now consider the Barkley system, and we show that the learning in 1-D can generalize to three spatial dimensions with Neumann (no-flux) boundary conditions on a cubic Cartesian domain $(x_{1},x_{2},x_{3})\in[-L,+L]\times[-L,+L]\times[-L,+L]\equiv\mathcal{C}$, i.e. for all fields $\Psi_{i}$.
$\star d\Psi_{i}\big|_{\partial\mathcal{C}}=0\ ,$ (27)
which is equivalent to stating that the normal derivative $\hat{e}_{\perp}\cdot\vec{\nabla}\Psi_{i}$ vanishes on $\partial\mathcal{C}$ (where $\hat{e}_{\perp}$ is the unit-vector normal to $\partial\mathcal{C}$ locally). These boundary conditions differ from those used during training and in previous cases.
We initialize the Barkley system with a field configuration that clearly displays scroll wave patterns (Appendix I.3). Fig. 3B1–4 shows that the neural network predictions, $\hat{U}(x_{1}(y_{1}),x_{2}(y_{2}),x_{3}(y_{3}),t)$ and $\hat{V}(x_{1}(y_{1}),x_{2}(y_{2}),x_{3}(y_{3}),t)$, closely match the true dynamics, $U(x_{1},x_{2},x_{3},t)$ and $V(x_{1},x_{2},x_{3},t)$, as governed by Eq. (5), demonstrating the adaptability of our dimension-free model to diverse boundary conditions.
D.3 Two-Dimensional Curved Surface Embedded in 3-Dimensional Euclidean Space
Figure 6: The time evolution of the FitzHugh–Nagumo system on a two-dimensional intrinsically curved surface. We illustrate the geographic and stereographic coordinate representations in (A–C), and in (D) compare the true dynamics obtained by integrating Eq. (22) in geographic coordinates with the learned dynamics generated in stereographic coordinates. In (A–B), we embed the spherical surface in three-dimensional Cartesian space $(x,y,z)$, showing the sphere in cyan, the equatorial plane in red, and their intersection—the equator—in purple: (A) shows a uniform grid in geographic coordinates $(\theta,\phi)$; (B) shows a uniform grid in stereographic coordinates $(\rho,\varphi)$—on the plane in (B1), projected onto the sphere in (B2), resulting in a non-uniform grid in $(\theta,\phi)$ as shown in (B3). (C1) shows a field $\Psi$ defined over $(\theta,\phi)$, and (C2) shows the same field over $(\rho,\varphi)$. In (D), we present field configurations at four time frames: (D1) $t=0$, (D2) $t=15$, (D3) $t=30$, (D4) $t=45$. Each time frame consists of four three-dimensional plots in embedded $(x,y,z)$-space: the top row shows the true dynamics, the bottom row shows the neural network predictions; the left column displays the $V$-field and the right column the $W$-field of the FitzHugh-Nagumo system.
In addition to increasing the number of the spatial dimensions and modifying the boundary conditions, we now bend the space (intrinsically). For this case, we explore the dynamics of the FitzHugh–Nagumo fields on a two-dimensional domain (similar to Fig. 3A but not flat) embedded in a spherical surface — a manifold with constant positive intrinsic curvature. We consider two coordinate representations of this curved space — the geographic $(\theta,\phi)$ and the stereographic $(\rho,\varphi)$ systems — as illustrated in Fig. 6A-C (See Appendix J for more details). The transformation between these coordinate systems is given by:
$\rho=R\tan\left(\phi/2\right)\ ,\ \varphi=\phi\ ,$ (28)
where $R$ is the radius of the sphere. These coordinates can be further related to the Cartesian coordinates $(x,y,z)$ of the ambient three-dimensional Euclidean embedding space via:
$x=R\sin\theta\sin\phi\ ,\ y=R\sin\theta\cos\phi\ ,\ z=R\cos\theta\ .$ (29)
The spatial domain will be constrained between the radial values $\rho_{\min}$ and $\rho_{\max}$, and the angular values $\varphi_{\min}$ and $\varphi_{\max}$, which can be determined from $\theta_{\min}$, $\theta_{\max}$, $\phi_{\min}$, $\phi_{\max}$ (Appendix I.4), We impose heterogeneous boundary conditions: the Neumann (no-flux) conditions on the latitude boundaries and the periodic conditions on the longitude boundaries.
We initialize the system with a clearly defined spiral pattern for the fields $V$ and $W$ (defined in $(\theta,\phi)$-coordinates, see Appendix I.4). The true dynamics — $V(x(\theta,\phi),y(\theta,\phi),z(\theta,\phi),t)$ and $W(x(\theta,\phi),y(\theta,\phi),z(\theta,\phi),t)$ — are obtained by integrating Eq.(4) in $(\theta,\phi)$-coordinates and then mapped to the embedding space $(x,y,z)$, as illustrated in the top row of Fig.6D1–4. For the stereographic $(\rho,\varphi)$-coordinates, we extrapolate the initial condition (Appendix H), integrate the neural network learned dynamics, and transform the results to $(x,y,z)$ to obtain $\hat{V}(x(\rho,\varphi),y(\rho,\varphi),z(\rho,\varphi),t)$ and $\hat{W}(x(\rho,\varphi),y(\rho,\varphi),z(\rho,\varphi),t)$, shown in the bottom row of Fig.6D1–4. We observe a remarkable agreement (quantified in Appendix K) between both predictions, illustrating that our machine learning method not only generalizes across coordinate systems, extends across dimensions, and adapts to different boundary conditions, but also remains robust under curvatures.
D.4 Two-Dimensional Flat Surface with Non-Trivial Topology
Finally, we change the topology of the environment. In essence, this is straightforward, since only the boundary specification needs to be modified. For example, let us study the FitzHugh–Nagumo system on a flat annulus in the polar coordinates $(\rho,\theta)$, which can be related to the Cartesian coordinates $(x,y)$ via:
$x=\rho\cos\theta\ ,\ y=\rho\sin\theta\ ,$ (30)
where the region of interests is confined within $\rho\in[\rho_{\min},\rho_{\max}]$, where we impose fixed-value Dirichlet boundary conditions.
Figure 7: The time evolution of the FitzHugh–Nagumo system on a two-dimension flat surface with an annulus topology. We present the time evolution of the true dynamics obtained by integrating Eq. (4) and compare it with the dynamics generated using the learned model at four different time frames: (A) $t=0$, (B) $t=15$, (C) $t=30$, (D) $t=45$. Each time frame consists of four two-dimensional plots: the top row shows the true dynamics, while the bottom row displays the neural network predictions; the left column corresponds to $V$-field and the right column to $W$-field of the FitzHugh-Nagumo system.
Similar to what has been done in Section D.2, we choose an initial field configuration at $t=0$ where the spiral wave pattern is clearly visible (an explanation of how we do it is provided in Appendix I.5). The true dynamics are computed by integrating Eq.(4) (Fig. 7A-D top row, $V(\rho,\theta,t)$ and $W(\rho,\theta,t)$). The learned model is integrated from the very same initial condition (Fig. 7A-D bottom row, $\hat{V}(\rho,\theta,t)$ and $\hat{W}(\rho,\theta,t)$). Both of these are mapped back to the $(x,y)$ coordinates, i.e. $V(x(\rho,\theta),y(\rho,\theta),t)$, $W(x(\rho,\theta),y(\rho,\theta),t)$, $\hat{V}(x(\rho,\theta),y(\rho,\theta),t)$, and $\hat{W}(x(\rho,\theta),y(\rho,\theta),t)$, for illustration purpose. We observe that the predictions agree, suggesting that our machine learning method not only generalizes across coordinate and intrinsically curved systems, but also applicable to non-trivial topologies. In other words, our “spatially liberated” approach is topology-free.
Appendix E Training the Neural Networks
We describe the training procedure and also the generation of training data within a one-dimensional periodic space equipped with a Cartesian coordinate where we impose periodic boundary conditions (Appendix B), i.e.
$x\in\mathcal{S}=[-L,+L]\ ,\ x\sim x+L\ ,$ (31)
and its use in training spatially liberated neural networks for each system as sketched in Fig. 1A.
For each of the systems (FitzHugh-Nagumo model, Barkley model, and Patlak-Keller-Segel model), we aim to train a neural network with an MLP architecture, consisting of $n_{lay}=2$ hidden layers, each with $n_{neu}$ neurons (to be specified). This network takes as input seven scalar quantities (0-forms) which are invariant features as listed in Eq. (1):
$\Psi_{1}\ ,\ \langle d\Psi_{1},d\Psi_{1}\rangle\ ,\ \star\ d\star d\Psi_{1}\ ,\ \Psi_{2}\ ,\ \langle d\Psi_{2},d\Psi_{2}\rangle\ ,\ \star\ d\star d\Psi_{2}\ ,\ \langle d\Psi_{1},d\Psi_{2}\rangle$ (32)
to predict two changing rates of the dynamical fields
$\partial_{t}\Psi_{1}\ ,\ \partial_{t}\Psi_{2}$
at each location $x$ in space and point $t$ in time; see Fig. 1A. For the FitzHugh-Nagumo system, we set $n_{\text{neu}}=64$, with the fields identified as $(\Psi_{1},\Psi_{2})=(V,W)$. In the Barkley system, we likewise use $n_{\text{neu}}=64$, where $(\Psi_{1},\Psi_{2})=(U,V)$. With the Patlak-Keller-Segel system, we consider $n_{\text{neu}}=32$, in which $(\Psi_{1},\Psi_{2})=(A,B)$.
The training data for each case is first simulated and then subsampled for training as described below: Section E.1 for the FitzHugh-Nagumo system, Section E.2 for the Barkley system, and Section E.3 for the modified Patlak-Keller-Segel system. For all simulations in this work, we have checked that the numerical computations are robust to space-time discretizations.
All computations in this work were performed on CPU-based machines. Each experiment was trained in under one hour using an internal compute node equipped with an Intel Core i7-7700 processor (3.60 GHz), 128 GB RAM, and local SSD storage. No GPUs or cloud services were used.
E.1 FitzHugh-Nagumo System Training
We generate numerical simulation data of the FitzHugh-Nagumo system on a one-dimensional domain specified in Eq. (31), using $L=10^{2}$. The spatiotemporal discretization is set to $(\Delta t,\Delta x)=(10^{-1},1)$, and for every numerical simulation of Eq. (22) we generate the time progression consecutively at each timestep $\Delta t$, using forward Euler method Eq. (21), for $t\in[0,T]$ where $T=10^{3}$. All training data is obtained from a single simulation, in which the initial field configuration — consisting of combinations of a few localized configurations — is given in Appendix F.1. Then, we randomly-selected a subset of the simulation data (i.e. $10\%$) to create the training dataset for neural network training. This reduces the training data size but still ensures a fine time-increment, minimizing error propagation and improving long-term prediction accuracy of the trained model. The neural network model is optimized using the least squares deviation criterion with the Adam optimizer Kingma (2014), set to a learning rate of $5\times 10^{-4}$, and trained for $2\times 10^{3}$ epochs.
E.2 Barkley System Training
We create numerical simulation data on the one-dimensional domain specified in Eq. (31) for Barkley system, using $L=5\times 10^{1}$. The spatiotemporal discretization is set to $(\Delta t,\Delta x)=(10^{-2},5\times 10^{-1})$, and for every numerical simulation of Eq. (23) we generate the time progression consecutively at each timestep $\Delta t$, using the forward Euler method Eq. (21), for $t\in[0,T]$ where $T=2\times 10^{2}$. All training data is obtained from a single simulation, in which the initial field configuration — consisting of combinations of a few localized configurations — is given in Appendix F.2. Then, we randomly-selected a subset of the simulation data (i.e. $5\%$) to create the training dataset for neural network training. This reduces the training data size but still ensures a fine time-increment, minimizing error propagation and improving long-term prediction accuracy of the trained model. The neural network model is optimized using the least squares deviation criterion with the Adam optimizer, set to a learning rate of $5\times 10^{-4}$, and trained for $2\times 10^{3}$ epochs.
E.3 Patlak-Keller-Segel System Training
We produce numerical simulation data for the Palak-Keller-Segel system on a one-dimensional domain specified in Eq. (31), using $L=10^{2}$. The spatiotemporal discretization is set to $(\Delta t,\Delta x)=(10^{-1},1$, and for every numerical simulation of Eq. (24) we generate the time progression consecutively at each timestep $\Delta t$, using forward Euler method Eq. (21), for $t\in[0,T]$ where $T=5\times 10^{2}$. The training data is derived from multiple simulations showcasing traveling wave collisions and diffusion. The initial field configurations used to create these simulations are uniform for $A$ and consist of combinations of a few localized configurations for $B$, as provided in Appendix F.3. Our choice of domain dimensions and initial field configurations are motivated by their relevance to real experimental settings using microfluidic devices. Then, we randomly-selected a subset of the simulation data (i.e. $2\%$) to create the training dataset for neural network training. This reduces the training data size but still ensures a fine time-increment, minimizing error propagation and improving long-term prediction accuracy of the trained model. The neural network model is optimized using the least squares deviation criterion with the Adam optimizer, set to a learning rate of $5\times 10^{-4}$, and trained for $2\times 10^{3}$ epochs.
For notational convenience, we redefine $A$, $B$ to represent the normalized quantities $A/A_{\text{max}}$, $B/B_{\text{max}}$ respectively. This convention applies from this point forward unless specified otherwise. Note that $A_{\max}=30$ and $B_{\max}=1$.
Appendix F Initial Field Configurations for Training
Here we list the initial field configurations used to generate the training data for the FitzHugh-Nagumo, the Barkley, and the Patlak-Keller-Segel models, as explained in Appendix E.
F.1 FitzHugh-Nagumo system
The initial field configurations are given by:
$\begin{split}V(x,0)&=V_{01}\exp\left[-\frac{(x-x_{V1})^{2}}{2\sigma_{xV1}}\right]+V_{02}\exp\left[-\frac{(x-x_{V2})^{2}}{2\sigma_{xV2}}\right]\ ,\\ W(x,0)&=W_{01}\exp\left[-\frac{(x-x_{W1})^{2}}{2\sigma_{xW1}}\right]\ ,\end{split}$ (33)
where:
$\begin{split}(V_{01},x_{V1},\sigma_{xV1})&=(+2,+40,10)\ ,\\ (V_{02},x_{V2},\sigma_{xV2})&=(-2,-40,10)\ ,\\ (W_{01},x_{W1},\sigma_{xW1})&=(+1,0,10)\ .\end{split}$ (34)
F.2 Barkley system
The initial field configurations are given by:
$\begin{split}U(x,0)&=U_{01}\exp\left[-\frac{(x-x_{U1})^{2}}{2\sigma_{xU1}}\right]\\ &+U_{02}\exp\left[-\frac{(x-x_{U2})^{2}}{2\sigma_{xU2}}\right]+U_{03}\exp\left[-\frac{(x-x_{U3})^{2}}{2\sigma_{xU3}}\right]\ ,\\ V(x,0)&=V_{01}\exp\left[-\frac{(x-x_{V1})^{2}}{2\sigma_{xV1}}\right]\\ &+V_{02}\exp\left[-\frac{(x-x_{V2})^{2}}{2\sigma_{xV2}}\right]+V_{03}\exp\left[-\frac{(x-x_{V3})^{2}}{2\sigma_{xV3}}\right]\ ,\end{split}$ (35)
where:
$\begin{split}(U_{01},x_{U1},\sigma_{xU1})&=(0.9,+22,4)\ ,\\ (U_{02},x_{U2},\sigma_{xU2})&=(0.5,-8,4)\ ,\\ (U_{03},x_{U3},\sigma_{xU3})&=(0.8,-28,2)\ ,\\ (V_{01},x_{V1},\sigma_{xV1})&=(0.5,0,4)\ ,\\ (V_{02},x_{V2},\sigma_{xV2})&=(0.9,20,2)\ ,\\ (V_{03},x_{V3},\sigma_{xV3})&=(0.7,-24,2)\ .\end{split}$ (36)
F.3 Patlak-Keller-Segel system
Here we consider an active matter system of many chemotactic bacteria that navigate in response to chemo-attractants, which they simultaneously deplete through their own consumption Phan et al. (2024). Two important phenomena emerge from bacterial collective chemotaxis: the formation of traveling bands of bacteria and the collisions of those bands, which then diffuse Phan et al. (2020). To generate a series of in silico experiments illustrating both phenomena, we begin with simple initial conditions: the Asp is uniform, and bacteria are inoculated at two ports, each modeled by a Gaussian distribution. We consider four different initial profiles for Asp concentration and four times three distinct initial profiles for the bacterial field, i.e.
$\begin{split}A(x,0)&=A_{0}\ ,\\ B(x,0)&=B_{1}\exp\left[-\frac{(x-x_{B1})^{2}}{2\sigma_{B1}}\right]+B_{2}\exp\left[-\frac{(x-x_{B2})^{2}}{2\sigma_{B2}}\right]\ ,\end{split}$ (37)
where:
$\begin{split}x_{B1}=-50\ ,\ x_{B2}=+50\ ,\ \sigma_{B1}=\sigma_{B2}=5\ &,\\ A_{0}\in A_{\max}\times\{0,1/4,1/2,3/4,1\}\ &,\\ B_{1}\in B_{\max}\times\{1/3,2/3,1\}\ ,\ B_{2}\in B_{\max}\times\{0,1/3,2/3,1\}\ &.\end{split}$ (38)
For notational convenience, we redefine $A$, $B$ to represent the normalized quantities $A/A_{\text{max}}$, $B/B_{\text{max}}$ respectively. This convention applies from this point forward unless specified otherwise. Note that $A_{\max}=30$ and $B_{\max}=1$. We have chosen the parameters so that these experiments are implementable within microfluidic environment.
Appendix G Cartesian and Non-Cartesian Coordinates
Here we show the relation between different choices of coordinate systems used in Section D.1 and Section D.2, from the mapping between locations to the calculation of relevant differential forms.
G.1 Inverse Function and Illustration of $y$ vs $x$
The inverse function, which helps calculating $y$ from $x$, is given by:
$y=f^{-1}(x)=\lambda\left[\frac{\Theta(x)}{\sqrt[3]{2}\cdot 3^{2/3}}-\frac{\sqrt[3]{\frac{2}{3}}}{\Theta(x)}\right]\ ,$ (39)
where
$\Theta(x)=\sqrt[3]{\sqrt{3}\sqrt{27\left(\frac{2x}{\lambda}\right)^{2}+4+9\left(\frac{2x}{\lambda}\right)}}\ .$
The transformations between Cartesian and non-Cartesian coordinates in different spaces are presented in Fig. 8.
Figure 8: The Cartesian and non-Cartesian coordinate systems in one- and two-dimensional space. Here we use Eq. (25) and Eq. (26) with $\lambda=L=10^{2}$. (A) One-dimensional space, conversion between $x$ and $y$. (B) Two-dimensional space, conversion between $(x_{1},x_{2})$ and $(y_{1},y_{2})$. The color transitions linearly with $x_{1}$, following the Python Matplotlib brg colormap Hunter (2007), shifting from blue to red to green.
G.2 Calculation of Differential-Forms in One-Dimensional Euclidean Space
Denote $^{\prime}=d/dy$ when taking the derivative of the function $f(y)$. The differential-form inputs (besides $\Psi_{1}$ and $\Psi_{2}$) used in the neural network as illustrated in Fig. 1B for the one-dimensional non-Cartesian case (Section D.1) can be calculated as follows:
  • The Laplacians:
    $\begin{split}\star d\star d\Psi_{j}(t)&=\frac{1}{\left|f^{\prime}(y)\right|}\left\{\frac{\left|f^{\prime}(y)\right|}{\left[f^{\prime}(y)\right]^{2}}\right\}^{\prime}\partial_{y}\Psi_{j}(y,t)+\frac{1}{\left[f^{\prime}(y)\right]^{2}}\partial_{y}^{2}\Psi_{j}(y,t)\ ,\end{split}$ (40)
    where $j=1,2$. There are two of such inputs.
  • The inner products:
    $\begin{split}\langle d\Psi_{j}(t),d\Psi_{k}(t)\rangle&=\left[\frac{1}{f^{\prime}(y)}\partial_{y}\Psi_{j}(y,t)\right]\left[\frac{1}{f^{\prime}(y)}\partial_{y}\Psi_{k}(y,t)\right]\ ,\end{split}$ (41)
    where $j,k=1,2$. There are three of such inputs, corresponding to three distinct inner products.
G.3 Calculation of Differential-Forms in Two-Dimensional Euclidean Space
The differential-form inputs (besides $\Psi_{1}$ and $\Psi_{2}$) used in the neural network as illustrated in Fig. 1B for the two-dimensional non-Cartesian case (Section D.2) can be calculated as follows:
  • The Laplacians:
    $\begin{split}\star d\star d\Psi_{j}(t)&=\frac{1}{\left|f^{\prime}(y_{1})f^{\prime}(y_{2})\right|^{1/2}}\partial_{y_{1}}\left\{\frac{\left|f^{\prime}(y_{1})f^{\prime}(y_{2})\right|^{1/2}}{\left[f^{\prime}(y_{1})\right]^{2}}\right\}\partial_{y_{1}}\Psi_{j}(y_{1},y_{2},t)\\ &+\frac{1}{\left|f^{\prime}(y_{1})f^{\prime}(y_{2})\right|^{1/2}}\partial_{y_{2}}\left\{\frac{\left|f^{\prime}(y_{1})f^{\prime}(y_{2})\right|^{1/2}}{\left[f^{\prime}(y_{2})\right]^{2}}\right\}\partial_{y_{2}}\Psi_{j}(y_{1},y_{2},t)\\ &+\frac{1}{\left[f^{\prime}(y_{1})\right]^{2}}\partial_{y_{1}}^{2}\Psi_{j}(y_{1},y_{2},t)+\frac{1}{\left[f^{\prime}(y_{2})\right]^{2}}\partial_{y_{2}}^{2}\Psi_{j}(y_{1},y_{2},t)\ ,\end{split}$ (42)
    where $j=1,2$. There are two of such inputs.
  • The inner products:
    $\begin{split}\langle d\Psi_{j}(t),d\Psi_{k}(t)\rangle=&\left[\frac{1}{f^{\prime}(y_{1})}\partial_{y_{1}}\Psi_{j}(y_{1},y_{2},t)\right]\left[\frac{1}{f^{\prime}(y_{1})}\partial_{y_{1}}\Psi_{k}(y_{1},y_{2},t)\right]\\ &+\left[\frac{1}{f^{\prime}(y_{2})}\partial_{y_{2}}\Psi_{j}(y_{1},y_{2},t)\right]\left[\frac{1}{f^{\prime}(y_{2})}\partial_{y_{2}}\Psi_{k}(y_{1},y_{2},t)\right]\ ,\end{split}$ (43)
    where $j,k=1,2$. There are three of such inputs, corresponding to three distinct inner products.
Appendix H Comparison between Different Coordinate Systems
We illustrate the comparison between coordinate systems — using 1D Cartesian and non-Cartesian coordinates as an example — which naturally extends to higher dimensions and curved spaces. For a given dynamical system, we initialize a condition (distinct from the training set) in $x$-space with uniform $\Delta x$ discretization. We then compute its evolution via both the true PDE and the neural network model from Section E, showing strong agreement. The corresponding initial condition in $y$-space (with uniform $\Delta y$) is evolved using the neural network, then mapped back to $x$-space via Eq. (25) for direct comparison. These steps are summarized in Fig. 9. For a coordinate-free model, all these generated datasets should be consistent, which we confirm through our results.
Figure 9: Testing coordinate-free learning in one-dimensional space. We study the time-evolution in different coordinates, using both “true” computed dynamics and ANN learned dynamics. (A) The physical domain of interest can be described (parametrized) by either the $x$-coordinate or the $y$-coordinate. An evenly-spaced $x$-grid is not equivalent to an evenly-spaced $y$-grid. (B) We start with an initial field configuration in $x$-space. (C1,C2) We extrapolate the values of the initial field configuration from what it looks like in $x$-space to its estimation in $y$-space. (D) In $x$-space, we compute the time-evolution of the field in two different ways: using the “true dynamics” (D1) and using the neural network learned dynamics (D2). We then can compare them. (E1) In $y$-space, we compute the time-evolution of the field using the neural network learned dynamics. (E2) We map the field back to physical space, showing what it looks like in $x$-space, then compare with (D).
Appendix I Initial Field Configurations for Testing
Here we list the initial field configurations used for testing in Section D.1 (see Appendix I.1), Section D.2 (see Appendix I.2), Section D.2 (see Appendix I.3), and Section D.3 (see Appendix I.4).
I.1 One-Dimensional Euclidean Space
In this case (Section D.1), we work with the Patlak-Keller-Segel system. The initial field configurations The initial field configurations used in testing follow similar structure to Eq. (38), with:
$\begin{split}x_{B1}=10\ ,\ x_{B2}=-60\ ,\ \sigma_{B1}=5\ ,\ \sigma_{B2}=6\ &,\\ A_{0}=5/6\ ,\ B_{1}=0.7\ ,\ B_{2}=0.4\ &.\end{split}$ (44)
We consider the same periodic space as in Appendix E for Patlak-Keller-Segel system, i.e. $x\in[-L,+L]$ with $L=10^{2}$ and a uniform grid of $\Delta x=1$. The transformation from this Cartesian coordinate $x$ to a non-Cartesian one $y$ follows Eq. (25). We use $y\in[-L,+L]$ with a uniform grid and a lattice spacing of $\Delta y=1$ when predicting the field evolution dynamics by integrating the trained neural network model from Section F.3 in this coordinate system.
We start with an initial field configuration that has not been used to generate the training data in Appendix E. The field $A$ is uniform everywhere and the field $B$ are combinations of localized configurations, as provided in Eq. (44). We create the field evolution dynamics using the timestep $\Delta t=10^{-1}$ for $t\in[0,5\times 10^{2}]$ by numerically integrating the PDEs Eq. (6).
I.2 Two-Dimensional Euclidean Space
In this case (Section D.2), we work with the FitzHugh-Nagumo system. We start with the field configuration given by:
$\begin{split}V(x_{1},x_{2})&=V_{01}\exp\left[-\frac{(x_{1}-{x_{1}}_{V1})^{2}}{2\sigma_{x1V1}}-\frac{(x_{2}-{x_{2}}_{V1})^{2}}{2\sigma_{x2V1}}\right]\\ &+V_{02}\exp\left[-\frac{(x_{1}-{x_{1}}_{V2})^{2}}{2\sigma_{x1V2}}-\frac{(x_{2}-{x_{2}}_{V2})^{2}}{2\sigma_{x2V2}}\right]\ ,\\ W(x_{1},x_{2})&=W_{01}\exp\left[-\frac{(x_{1}-{x_{1}}_{W1})^{2}}{2\sigma_{x1W1}}-\frac{(x_{2}-{x_{2}}_{W1})^{2}}{2\sigma_{x2W1}}\right]\ ,\end{split}$ (45)
where:
$\begin{split}(V_{01},{x_{1}}_{V1},{x_{2}}_{V1},\sigma_{x1V1},\sigma_{x2V1})&=(+1,0,0,15,10)\ ,\\ (V_{02},{x_{1}}_{V2},{x_{2}}_{V2},\sigma_{x1V2},\sigma_{x2V2})&=(+1,0,+40,8,8)\ ,\\ (W_{01},{x_{1}}_{W1},{x_{2}}_{W1},\sigma_{x1W1},\sigma_{x2W1})&=(+1,+10,0,10,25)\ .\end{split}$ (46)
Both $V$ and $W$ are combinations of localized configurations. Then, we numerically integrate the PDEs Eq. (4) for the total amount of time $\Delta T=8\times 10^{2}$ (so that the system has reached the stage of steady spiral wave generation) from the above configurations. We use the final-time configuration as the initial field configuration for the study in Section D.2.
Consider the periodic spatial domain $(x_{1},x_{2})\in[-L,+L]\times[-L,+L]$ with $L=10^{2}$ and uniform grid of lattice-spacing $(\Delta x_{1},\Delta x_{2})=(1,1)$. The transformation from these Cartesian coordinates $(x_{1},x_{2})$ to non-Cartesian ones $(y_{1},y_{2})$ is as described Eq. (26). We use $(y_{1},y_{2})\in[-L,+L]\times[-L,+L]$ with a uniform grid and a lattice-spacing of $(\Delta y_{1},\Delta y_{2})=1$ to describe the same space in this new coordinate system.
Since $(x_{1},x_{2})$-grid and $(y_{1},y_{2})$-grid do not represent identical set of physical points, we need to extrapolate the configuration Eq. (46) to $(y_{1},y_{2})$-coordinate. We generate the field evolution dynamics using the timestep $\Delta t=10^{-1}$ for $t\in[0,10^{2}]$, in different coordinate systems.
I.3 Three-Dimensional Euclidean Space
In this case (Section D.2), we work with the Barkley system. We start with the field configuration given by:
$\begin{split}U(x_{1},x_{2},x_{2})&=U_{01}\exp\left[-\frac{(x_{1}-{x_{1}}_{U1})^{2}}{2\sigma_{x1U1}}-\frac{(x_{2}-{x_{2}}_{U1})^{2}}{2\sigma_{x2U1}}-\frac{(x_{3}-{x_{3}}_{U1})^{2}}{2\sigma_{x3U1}}\right]\\ &+U_{02}\exp\left[-\frac{(x_{1}-{x_{1}}_{U2})^{2}}{2\sigma_{x1U2}}-\frac{(x_{2}-{x_{2}}_{U2})^{2}}{2\sigma_{x2U2}}-\frac{(x_{3}-{x_{3}}_{U2})^{2}}{2\sigma_{x3U2}}\right]\ ,\\ V(x_{1},x_{2},x_{3})&=V_{01}\exp\left[-\frac{(x_{1}-{x_{1}}_{V1})^{2}}{2\sigma_{x1V1}}-\frac{(x_{2}-{x_{2}}_{V1})^{2}}{2\sigma_{x2V1}}-\frac{(x_{3}-{x_{3}}_{V1})^{2}}{2\sigma_{x3V1}}\right]\\ &+V_{02}\exp\left[-\frac{(x_{1}-{x_{1}}_{V2})^{2}}{2\sigma_{x1V2}}-\frac{(x_{2}-{x_{2}}_{V2})^{2}}{2\sigma_{x2V2}}-\frac{(x_{3}-{x_{3}}_{V2})^{2}}{2\sigma_{x3V2}}\right]\ ,\end{split}$ (47)
where:
$\begin{split}(U_{01},{x_{1}}_{U1},{x_{2}}_{U1},{x_{3}}_{U1},\sigma_{x1U1},\sigma_{x2U1},\sigma_{x3U1})&=(0.5,0,6,0,5,6,5)\ ,\\ (U_{02},{x_{1}}_{U2},{x_{2}}_{U2},{x_{3}}_{U2},\sigma_{x1U2},\sigma_{x2U2},\sigma_{x3U2})&=(0.5,0,12,6,4,5,4)\ ,\\ (V_{01},{x_{1}}_{V1},{x_{2}}_{V1},{x_{3}}_{V1},\sigma_{x1V1},\sigma_{x2V1},\sigma_{x3V1})&=(0.8,5,6,0,5,4,4)\ ,\\ (V_{02},{x_{1}}_{V2},{x_{2}}_{V2},{x_{3}}_{V2},\sigma_{x1V2},\sigma_{x2V2},\sigma_{x3V2})&=(0.7,-7,6,-7,4,4,6)\ .\end{split}$ (48)
Both $U$ and $V$ are combinations of localized configurations. Then, we numerically integrate the PDEs Eq. (5) for the total amount of time $\Delta T=2\times 10^{1}$ (so that the system has reached the stage of steady scroll wave generation) from the above configurations. We use the final-time configuration as the initial field configuration for the study in Section D.2.
We consider the confined space $(x_{1},x_{2},x_{3})\in[-L,+L]\times[-L,+L]\times[-L,+L]$ with $L=5\times 10^{1}$, covered by a uniform grid of lattice-spacing $(\Delta x_{1},\Delta x_{2},\Delta x_{3})=(5\times 10^{-1},5\times 10^{-1},5\times 10^{-1})$. We compute the field evolution dynamics in this Cartesian coordinate system using the timestep $\Delta t=10^{-2}$ for $t\in[0,5]$.
I.4 A Two-Dimensional, Intrinsically Curved Surface
In this case (Section D.3), we work with the FitzHugh-Nagumo system. On the spherical surface, the FitzHugh-Nagumo system written in the $(\theta,\phi)$-coordinate can be described by the following PDEs:
$\begin{split}\partial_{t}V(\theta,\phi,t)&=\frac{1}{R^{2}}\left(\cot\theta\partial_{\theta}+\partial_{\theta}^{2}+\frac{1}{\sin^{2}\theta}\partial_{\phi}^{2}\right)V(\theta,\phi,t)\\ &\ \ \ \ \ \ \ \ +V(\theta,\phi,t)-\frac{1}{3}V^{3}(\theta,\phi,t)-W(\theta,\phi,t)\ ,\\ \partial_{t}W(\theta,\phi,t)&=\varepsilon\left[V(\theta,\phi,t)+\beta-\gamma W(\theta,\phi,t)\right]\ .\end{split}$ (49)
We start with the field configuration given by:
$\begin{split}V(\rho,\varphi,0)&=V_{01}\exp\left[-\frac{(\rho-\rho_{V1})^{2}}{2\sigma_{\rho V1}}-\frac{(\varphi-\varphi_{V1})^{2}}{2\sigma_{\varphi V1}}\right]\\ &+V_{02}\exp\left[-\frac{(\rho-\rho_{V2})^{2}}{2\sigma_{\rho V2}}-\frac{(\varphi-\varphi_{V2})^{2}}{2\sigma_{\varphi V2}}\right]\ ,\\ W(\rho,\varphi,0)&=W_{0}\exp\left[-\frac{(\rho-\rho_{W})^{2}}{2\sigma_{\rho W}}-\frac{(\varphi-\varphi_{W})^{2}}{2\sigma_{\varphi W}}\right]\ .\end{split}$ (50)
where:
$\begin{split}(V_{01},\rho_{V1},\varphi_{V1},\sigma_{\rho V1},\sigma_{\varphi V1})&=(1,140,\pi/6,15,0.15)\ ,\\ (V_{02},\rho_{V2},\varphi_{V2},\sigma_{\rho V2},\sigma_{\varphi V2})&=(1,160,\pi/4,10,0.1)\ ,\\ (W_{0},\rho_{W},\varphi_{W},\sigma_{\rho W},\sigma_{\varphi W})&=(1,150,\pi/4,10,0.2)\ .\end{split}$ (51)
Both $V$ and $W$ are combinations of localized configurations. Then, we numerically integrate the PDEs Eq. (49) for the total amount of time $\Delta T=2\times 10^{2}$ (so that the system has reached the stage of steady spiral wave generation) from the above configurations. We use the final-time configuration as the initial field configuration for the study in Section D.3.
We work on the spherical surface of radius $R=2\times 10^{2}$, analyzing the two-dimensional curved space covered by $(\rho,\varphi)\in[\rho_{\min},\rho_{\max}]\times[\varphi_{\min},\varphi_{\max}]$, where $\rho_{\min}=5\times 10^{1}$, $\rho_{\max}=2\times 10^{2}$ and $\varphi_{\min}=0$, $\varphi_{\max}=\pi/2$. We discretize the coordinates $(\rho,\varphi)$ uniformly with the lattice-spacing
$\left(\Delta\rho,\Delta\varphi\right)=\left(\frac{\rho_{\max}-\rho_{\min}}{300},\frac{\varphi_{\max}-\varphi_{\min}}{200}\right)\ .$ (52)
The transformation from these stereographic coordinates $(\rho,\varphi)$ to geographical ones $(\theta,\phi)$ is as described Eq. (28). We use $(\theta,\phi)\in[-L,+L]\times[\theta_{\min},\theta_{\max}]\times[\phi_{\min},\phi_{\max}]$ to describe the same space in this new coordinate system, where we use a uniform grid with the lattice-spacing
$\left(\Delta\theta,\Delta\phi\right)=\left(\frac{\theta_{\max}-\theta_{\min}}{300},\frac{\phi_{\max}-\phi_{\min}}{200}\right)\ .$ (53)
Note that the $(\rho,\varphi)$-grid and the $(\theta,\phi)$-grids are not comprised of identical sets of physical points. We compute the field evolution dynamics on the spherical surface in this different coordinate system, using the timestep $\Delta t=10^{-1}$ for $t\in[0,5\times 10^{1}]$.
I.5 A Two-Dimensional, Flat Surface, with Non-Trivial Topology
We work with the FitzHugh-Nagumo system for this demonstration (Section D.4). In 2-D polar coordinate system, the FitzHugh-Nagumo PDEs can be expressed in the $(\rho,\theta)$-coordinate as:
$\begin{split}\partial_{t}V(\rho,\theta,t)&=\frac{1}{\rho}\partial_{\rho}\left[\rho\partial_{\rho}V(\rho,\theta,t)\right]+\frac{1}{\rho^{2}}\partial_{\theta}^{2}V(\rho,\theta,t)\\ &\ \ \ \ \ \ \ \ +V(\rho,\theta,t)-\frac{1}{3}V^{3}(\rho,\theta,t)-W(\rho,\theta,t)\ ,\\ \partial_{t}W(\rho,\theta,t)&=\varepsilon\left[V(\theta,\phi,t)+\beta-\gamma W(\theta,\phi,t)\right]\ .\end{split}$ (54)
We start with the field configuration given by:
$\begin{split}V(\rho,\theta,0)&=V_{01}\exp\left[-\frac{(\rho-\rho_{V1})^{2}}{2\sigma_{\rho V1}}-\frac{(\theta-\theta_{V1})^{2}}{2\sigma_{\theta V1}}\right]\\ &+V_{02}\exp\left[-\frac{(\rho-\rho_{V2})^{2}}{2\sigma_{\rho V2}}-\frac{(\theta-\theta_{V2})^{2}}{2\sigma_{\theta V2}}\right]\\ &+V_{01}\exp\left[-\frac{(\rho-\rho_{V1})^{2}}{2\sigma_{\rho V1}}-\frac{(\theta-\theta_{V1}-\pi/2)^{2}}{2\sigma_{\theta V1}}\right]\\ &+V_{02}\exp\left[-\frac{(\rho-\rho_{V2})^{2}}{2\sigma_{\rho V2}}-\frac{(\theta-\theta_{V2}-3\pi/2)^{2}}{2\sigma_{\theta V2}}\right]\ ,\\ W(\rho,\varphi,0)&=W_{0}\exp\left[-\frac{(\rho-\rho_{W})^{2}}{2\sigma_{\rho W}}-\frac{(\theta-\theta_{W})^{2}}{2\sigma_{\theta W}}\right]\\ &+W_{0}\exp\left[-\frac{(\rho-\rho_{W})^{2}}{2\sigma_{\rho W}}-\frac{(\theta-\theta_{W}-\pi/2)^{2}}{2\sigma_{\theta W}}\right]\\ &+W_{0}\exp\left[-\frac{(\rho-\rho_{W})^{2}}{2\sigma_{\rho W}}-\frac{(\theta-\theta_{W}-3\pi/2)^{2}}{2\sigma_{\theta W}}\right]\ .\end{split}$ (55)
where:
$\begin{split}(V_{01},\rho_{V1},\theta_{V1},\sigma_{\rho V1},\sigma_{\theta V1})&=(1,140,\pi/6,15,0.15)\ ,\\ (V_{02},\rho_{V2},\theta_{V2},\sigma_{\rho V2},\sigma_{\theta V2})&=(1,160,\pi/4,10,0.1)\ ,\\ (W_{0},\rho_{W},\theta_{W},\sigma_{\rho W},\sigma_{\theta W})&=(1,150,\pi/4,10,0.2)\ .\end{split}$ (56)
Both $V$ and $W$ are combinations of localized configurations. Then, we numerically integrate the PDEs Eq. (54) for the total amount of time $\Delta T=10^{3}$ (so that the system has reached the stage of steady spiral wave generation) from the above configurations. We use the final-time configuration as the initial field configuration for the study in Section D.4.
We work on the annulus confined between the radius value $\rho\in[\rho_{\min},\rho_{\max}]$, where $\rho_{\min}=10^{2}$, $\rho_{\max}=2\times 10^{2}$. We discretize the coordinates $(\rho,\theta)$ uniformly with the lattice-spacing
$\left(\Delta\rho,\Delta\theta\right)=\left(1,2\pi/800\right)\ .$ (57)
Then, we compute the field evolution dynamics on the spherical surface with the true dynamics and with the neural network learned dynamics, using the timestep $\Delta t=10^{-1}$ for $t\in[0,5\times 10^{1}]$.
Appendix J Geographic and Stereographic Coordinates
We consider a sphere of radius $R$, equipped with geographical-coordinates $(\theta,\phi)$ corresponding to spherical-angles, where $\theta=0$ represents the South pole and $\theta=\pi$ is the North pole. The spherical surface can be embedded in a three-dimensional Euclidean space via Eq. (29), where $(x,y,z)$ are the corresponding Cartesian coordinates. On this curved manifold, the induced metric is given by:
$ds^{2}=R^{2}\left(d\theta^{2}+\sin^{2}\theta d\phi^{2}\right)\ .$ (58)
There, Laplacians and the inner-products can be calculated from:
  • The Laplacians:
    $\begin{split}\star d\star d\Psi_{j}&=\frac{1}{R^{2}}\left(\cot\theta\partial_{\theta}\Psi_{j}+\partial_{\theta}^{2}\Psi_{j}+\frac{1}{\sin^{2}\theta}\partial_{\phi}^{2}\Psi_{j}\right)\ ,\end{split}$ (59)
    where $j=1,2$. There are two of them.
  • The inner products:
    $\begin{split}\langle d\Psi_{j},d\Psi_{k}\rangle&=\frac{1}{R^{2}}\left(\partial_{\theta}\Psi_{j}\partial_{\theta}\Psi_{k}+\frac{1}{\sin^{2}\theta}\partial_{\phi}\Psi_{j}\partial_{\phi}\Psi_{k}\right)\ ,\end{split}$ (60)
    where $j,k=1,2$. There are three distinct inner products.
We study a region of the sphere surface bounded by latitude $\theta_{\min}$ and $\theta_{\max}$, and by longitude $\phi_{\min}$ and $\phi_{\max}$. We also impose the Neumann (no-flux) conditions as described in Eq. (27) on the latitude boundaries and the periodic conditions on the longitude boundaries.
We focus on a projection that utilizes the South pole as the projection viewpoint and the equator plane as the projection plane. A convenient mapping — written in stereographic polar-coordinates — is as introduced in Eq. (28). Thus, the metric in Eq. (58) after this coordinate transformation becomes:
$ds^{2}=\left[\frac{4R^{4}}{(R^{2}+\rho^{2})^{2}}\right]\left(d\rho^{2}+\rho^{2}d\varphi^{2}\right)\ .$ (61)
The Laplacians and the inner-products in these new coordinates can be expressed with:
  • The Laplacians:
    $\begin{split}\star d\star d\Psi_{j}&=\left[\frac{(R^{2}+\rho^{2})^{2}}{4R^{4}}\right]\left(\frac{1}{\rho}\partial_{\rho}\Psi_{j}+\partial_{\rho}^{2}\Psi_{j}+\frac{1}{\rho^{2}}\partial_{\varphi}^{2}\Psi_{j}\right)\ ,\end{split}$ (62)
    where $j=1,2$. There are two of them.
  • The inner products:
    $\begin{split}\langle d\Psi_{j},d\Psi_{k}\rangle&=\left[\frac{(R^{2}+\rho^{2})^{2}}{4R^{4}}\right]\left(\partial_{\rho}\Psi_{j}\partial_{\rho}\Psi_{k}+\frac{1}{\rho^{2}}\partial_{\rho}\Psi_{j}\partial_{\rho}\Psi_{k}\right)\ ,\end{split}$ (63)
    where $j,k=1,2$. There are three distinct inner products.
In this coordinate system, the spatial domain of interests is bound between radius $\rho_{\min}$ and $\rho_{\max}$, and angle $\varphi_{\min}$ and $\varphi_{\max}$, which can be determined from $\theta_{\min}$, $\theta_{\max}$, $\phi_{\min}$, $\phi_{\max}$, with Eq. (28). Figure 6A-C illustrates how geographical coordinates $(\theta,\phi)$ and stereographic coordinates $(\rho,\varphi)$ are related through a projection from the South pole onto the equatorial plane. It shows that the same region on the surface of the sphere can be covered by evenly-spacing grids in both coordinate systems, even though these grids are not equivalent.
We aim to demonstrate that, using the neural network model trained in one dimension in Sec. E, we can predict the correct field evolution dynamics on the surface of the sphere regardless of the chosen coordinate system. In other words, not only do we need to generate a prediction, but also compare it with the expected true dynamics. We begin with an initial condition defined in $(\theta,\phi)$-space with uniform $(\Delta\theta,\Delta\phi)$ discretization. We then compute the field evolution in the $(\theta,\phi)$-coordinate using direct integration of the true PDE. Next, we determine by extrapolation the corresponding initial condition in $(\rho,\varphi)$-space with uniform $(\Delta\rho,\Delta\varphi)$ discretization and use it to generate the field evolution in $(\rho,\varphi)$-coordinates via the trained neural network model in Section E. To make comparison, we transform both the generated datasets to the three-dimensional Euclidean $(x,y,z)$-space. These steps are similar to what has been demonstrated for one-dimensional case, see Fig. 9. For our spatially liberated neural network model, all these generated datasets are expected to be in agreement, which we confirm through our results.
Appendix K Comparing “True” and Learned Dynamics
To compare the “true” dynamics $\Psi_{i}(\zeta,t)$ with the neural learned dynamics $\hat{\Psi}(\eta,t)$ (which computed at different locations), we extrapolate from $\hat{\Psi}$ evaluated at $\eta$-locations to its corresponding values at $\zeta$-locations, denoted as $\tilde{\Psi}_{i}(\eta,t)$. We then can calculate $\Psi_{i}-\tilde{\Psi}_{i}$ to see the deviation between these dynamics and report the mean square-error (MSE) for all cases.
K.1 1-, 2- and 3-Dimensional Euclidean Space
We show the deviation for Patlak-Keller-Segel system in one-dimensional Euclidean space (Fig. 10A), FitzHugh-Nagumo system in two-dimensional Euclidean space (Fig. 10B), and Barkley system in three-dimensional Euclidean space (Fig. 10C).
[…]
  • Fig. 10A1: $_{A}=1.67\times 10^{-5}, MSE$$_{B}=4.90\times 10^{-6}$.
  • Fig. 10A2: $_{A}=8.56\times 10^{-6}, MSE$$_{B}=7.76\times 10^{-5}$.
  • Fig. 10A3: $_{A}=1.24\times 10^{-5}, MSE$$_{B}=7.98\times 10^{-5}$.
  • Fig. 10B1: $_{V}=2.01\times 10^{-3}, MSE$$_{W}=1.20\times 10^{-5}$.
  • Fig. 10B2: $_{V}=1.62\times 10^{-2}, MSE$$_{W}=1.49\times 10^{-4}$.
  • Fig. 10B3: $_{V}=3.03\times 10^{-2}, MSE$$_{W}=9.16\times 10^{-4}$.
  • Fig. 10B4: $_{V}=4.67\times 10^{-2}, MSE$$_{W}=1.39\times 10^{-3}$.
  • Fig. 10C1: $_{U}=0, MSE$$_{V}=0$ (initial field configurations are identical).
  • Fig. 10C2: $_{U}=3.19\times 10^{-3}, MSE$$_{V}=3.35\times 10^{-4}$.
  • Fig. 10C3: $_{U}=6.24\times 10^{-3}, MSE$$_{V}=7.44\times 10^{-4}$.
  • Fig. 10C4: $_{U}=9.31\times 10^{-3}, MSE$$_{V}=1.15\times 10^{-3}$.
Figure 10: Comparisons between “true” and neural learned dynamics in Euclidean spaces. The deviation range for each field is equal to $10\%$ its dynamical range, with negative deviation colored red-ish and positive deviation colored blue-ish. For Patlak-Keller-Segel system, we show the deviation between Fig. 2B1 and Fig. 2B2 in (A1), between Fig. 2C and Fig. 2D1 in (A2), and between Fig. 2B1 and Fig. 2D2 in (A3). For FitzHugh-Nagumo system, we show the deviation between the top and bottom-rows in Fig. 3A1-4 in (B1-4), respectively. For FitzHugh-Nagumo system, we show the deviation between the top and bottom-rows in Fig. 3B1-4 in (C1-4), respectively. Note that, for brevity, we only label the first plot in (B) and (C), since the rest are identical.
K.2 2-Dimensional Intrinsically curved Surface
We show the deviation for FitzHugh-Nagumo system in two-dimensional curved space, which is embedded in three-dimensional Euclidean space, in Fig. 11.
[…]
  • Fig. 11A: $_{V}=5.85\times 10^{-5}, MSE$$_{W}=1.35\times 10^{-8}$.
  • Fig. 11B: $_{V}=1.36\times 10^{-3}, MSE$$_{W}=1.51\times 10^{-5}$.
  • Fig. 11C: $_{V}=2.34\times 10^{-3}, MSE$$_{W}=6.56\times 10^{-5}$.
  • Fig. 11D: $_{V}=4.47\times 10^{-3}, MSE$$_{W}=1.01\times 10^{-4}$.
Figure 11: Comparisons between “true” and neural learned dynamics in two-dimensional intrinsically curved surface embedded in three-dimensional Euclidean space. Here we study FitzHugh-Nagumo system, where the deviation range for each field is equal to $10\%$ of its dynamical range, with negative deviation colored red-ish and positive deviation colored blue-ish. We show the deviation between the top and bottom rows of Fig. 6D1 in (A), of Fig. 6D2 in (B), of Fig. 6D3 in (C), and of Fig. 6D4 in (D). Note that, for brevity, we only label the first plot, since the rest are identical.
K.3 2-Dimensional Non-Trivial Topology
We show the deviation for FitzHugh-Nagumo system in two-dimensional flat space with an annulus topology, in Fig. 12.
[…]
  • Fig. 12A: $_{V}=0, MSE$$_{W}=0$ (initial field configurations are identical).
  • Fig. 12B: $_{V}=9.25\times 10^{-4}, MSE$$_{W}=1.11\times 10^{-5}$.
  • Fig. 12C: $_{V}=1.64\times 10^{-3}, MSE$$_{W}=4.69\times 10^{-5}$.
  • Fig. 12D: $_{V}=3.33\times 10^{-3}, MSE$$_{W}=7.54\times 10^{-5}$.
Figure 12: Comparisons between “true” and neural learned dynamics in two-dimensional flat surface embedded with an annulus topology. Here we study FitzHugh-Nagumo system, where the deviation range for each field is equal to $10\%$ of its dynamical range, with negative deviation colored red-ish and positive deviation colored blue-ish. We show the deviation between the top and bottom rows of Fig. 7A in (A), of Fig. 7B in (B), of Fig. 7C in (C), and of Fig. 7D in (D).
Appendix L On the Quality of the Prediction using Noisy Training Data
In fact, our study was developed with real experiments in mind: we are preparing to apply it to a real bacteria-chemical system as reported in Phan et al. (2024), which is also the work that proposed the modified Patlak-Keller-Segel PDE used in the main text. The model in Eq. (24) is presented in dimensionless form; Table 1 summarizes the mapping between dimensional parameters and their dimensionless counterparts.
Asp dynamics Bacteria dynamics Non-Dimensionalized
[…]
Table 1: The best-estimated parameters are those reported in Phan et al. (2024), based on an analysis of traveling-wave experiments with E. coli strain HE205 Cremer et al. (2019). Unlisted non-dimensionalized parameters are, by convention, implied to be set to $1$, i.e. $D_{A}=1$, $\gamma_{0}=1$, $B_{C}=1$, and $A_{h}=1$. In those experiments, the maximum Asp concentration $\approx 100\mu\mathrm{M}$ and the maximum cell density $\approx 6\mathrm{OD}$, which correspond to the nondimensional maxima $A_{\max}\approx 27$ and $B_{\max}\approx 1$. The non-dimensionalized unit-time and unit-length correspond to the time-scale $\tau\approx 1.4\mathrm{s}$ and the length-scale $\lambda\approx 33\mu\mathrm{m}$. While these values can vary across strains, they are expected to remain within the same order of magnitude.
We can show that the experimental noise for this system is relatively low, can be made much lower, and also can be estimated as follows. The microfluidic channel has width $W\approx 1.2\mathrm{mm}$ and depth $H\approx 100\mu\mathrm{m}$ (which can be increased). The bacterial density in the traveling wave is around $n\sim 2\mathrm{OD}$ (with $1\mathrm{OD}\approx 8\times 10^{8}\mathrm{cells/mL}$), and wave peaks can reach $\sim 5\mathrm{OD}$. The effective one-dimensional bin size in our discretization is $\Delta x=1$ (dimensionless), corresponding to $\approx 33\mu\mathrm{m}$ physically (see the caption of Table 1). Hence, the expected number of bacteria per bin is
$N\sim nWH\Delta x\approx 8\times 10^{3}\mathrm{cells}\ .$
Assuming purely Poisson (counting) statistics, which approach a Gaussian in the large-number limit, the relative shot noise in the bacterial density field $B(x,t)$ is $\sigma\sim N^{-1/2}\sim 1\%$. Using a deeper device, e.g. increasing $H$ to $1.2\mathrm{mm}$, would reduce this further to
$\sigma\approx 0.3\%\ .$
The chemical concentration field $A(x,t)$ is even less noisy (effectively negligible in experiments), owing to its much higher molecular counts and diffusivity compared to those of bacteria. We believe these empirical conditions make the noise-free assumption acceptable here, while still laying groundwork for further studies.
If we add this $\sigma\sim 0.3\%$ noise to the training data and apply Savitzky-Golay smoothing before computing derivatives, the neural network – after the same training mentioned in Appendix E.3 – can still make reasonable and correct-qualitative predictions (especially in the region where both bacteria and chemical field values are high, where noise does not dominate training, which is from $t=0$ to $250$) for the initial condition as used in Fig. 2B1 and Fig. 2B2 of the main manuscript. Here we report the MSE for the values between $t=0$ and $t=250$ (see Fig. 13A):
  • $_{A}=4.76\times 10^{-3}, MSE$$_{B}=1.24\times 10^{-3}$.
Despite a substantially higher MSE after adding noise, the practical implication does not change: the prediction correctly tracks the traveling-wave front (speed, shape, persistence). Also, it should be noted that our current noise‐handling strategy is simple – simply applying a mild smoothing filter before training – but already yields acceptable predictions.
Figure 13: The predictions for Patlak-Keller-Segel field evolution dynamics in a periodic one-dimensional space, using training data corrupted by realistic noise. The initial field configurations are the same as used in Fig. 2B1 and Fig. 2B2. The gray boxes capture the spacetime regions between $t=0$ and $t=200$. (A1) The time progression of the fields when using the neural network trained the same way as in Appendix E.3. (A2) The field values difference between the neural network predicted Fig. 13A1 and the true Fig. 2B1. (B1) The time progression of the fields when using the neural network trained with $L^{1}$-norm (instead of $L^{2}$-norm). (B2) The field values difference between the neural network predicted Fig. 13B1 and the true Fig. 2B1.
There are several ways to improve performance; for instance, changing the training procedure by switching from an $L^{2}$-norm to a $L^{1}$-norm for the loss function makes the neural network model much more robust to noise. For comparison, we report this MSE between $t=0$ and $t=250$:
  • $_{A}=5.26\times 10^{-4}, MSE$$_{B}=1.32\times 10^{-4}$.
These results are now smaller by more than an order of magnitude and could be reduced further with more sophisticated noise‐resistant strategies.
Appendix M General PDE Example
The PCA model was fit to the 200 numerical trajectories of the 1-D harmonic oscillator. For these solutions, we approximate the first and second derivatives of $u$ using finite differences on a set of 200 solutions (with random initial conditions), and perform PCA, we obtained explained variance ratios of 6.7e1,3.3e1,6.4e66.7e-1,3.3e-1,6.4e-6, denoting the 2-dimensional linear structure of the manifold.
After identifying the operator, we proceed to train a PINN in the manner discussed in the main text. The PINN architecture consists of three layers with width 64 and a $\tanh$ activation function, trained using Adam Kingma (2014). The mean square error between a classical PINN solution and the coordinate-free version visualized in Fig. 14 is approximately $4e^{-4}$, while the square difference between the these solutions is depicted the right sub-plot of Figure 14.
Figure 14: Left: Solution of the Helmholtz equation obtained using classical PINN training. Middle: Solution of the Helmholtz equation obtained using a PINN with a coordinate-free PDE loss, where the PDE operator is identified from trajectories of the 1-D harmonic oscillator (see text for details). Right: Squared error between classical PINN solution and coordinate-free PINN solution to the 2D helmholtz problem.
Appendix N Vector PDE Example
The solution to the flow around a cylinder example is computed on the unit square, $(x,y)\in[0,1]^{2}\subset\mathbb{R}^{2}$, discretized using a regular grid of $100\times 100$ points. A cylinder of radius $r=0.1$ is centered at the center of the domain. The velocity field $v(x,y,t)$ is set initially to zero $v(x,y,0)=(0,0)$, with the exception of the left boundary of the square where we set $v(0,x,t)=1$. The viscosity $\nu=\eta/\rho$ is set to $\nu=0.01$.
After obtaining a numerical solution $\hat{v}$ using the setup described above, we compute the corresponding vector features $\quantity{(v\cdot\gradient)v,\Delta v,\gradient p}$. Using least squares, we determine the coefficients that best approximate $\partialderivative{v}{t}$ as a linear combination of these features, which for the example presented in Figure 5 are $(1.01,0.01,1.00)$ (rounded to decimal digits).
Using these coefficients, we integrate the momentum equation for the Taylor-Green vortex problem described below, using an Euler integration scheme, while assuming that the true pressure field is known for all time. We integrate on the domain $0,2\pi]^{2}\subset\mathbb{R}^{2}$, discretized using a $100\times 100$ grid. The solution depicted in the bottom of Figure 5 is obtained after 10 Euler time steps of size $dt=0.01$, with a relative $\ell^{2}$ error of 1.2e-1 (with respect to the analytical solution).
The Taylor-Green Vortex dynamics (Taylor & Green, 1937) are a special case of the incompressible Navier-Stokes equation, defined as follows:
For the (periodic) domain ${(x,y)}\in[0,2\pi]^{2}$, we consider the initial velocity field to take the form
$v_{1}=\cos(x)\sin(y),\quad v_{2}=-\sin(x)\cos(y)$ (64)
In this case, it is known that the velocity field as a function of time takes the form:
$v(x,y,t)$ $=v(x,y,0)e^{-2\frac{\eta}{\rho}t}$ (65)
$p(x,y,t)$ $=\frac{\rho}{4}\quantity({\cos(2x)+\cos(2y)})e^{-4\frac{\eta}{\rho}t}$ (66)
To further simplify the PDE identification and integration problem, we assume that the scalar field $p(x,y,t)$ is known at all times, and that the continuity equation is also known and does not need to be identified. In that way, we have reduced the problem to identifying a time-evolution vector-PDE (i.e. the momentum equation) for which we can use a simple Euler scheme to integrate (in order to verify our coordinate-free learning procedure).